### v1

# First Comparative Exoplanetology Within a Transiting Multi-planet System: Comparing the atmospheres of V1298 Tau b and c

11email: s.barat@uva.nl 22institutetext: School of Earth and Planetary Sciences (SEPS), National Institute of Science Education and Research (NISER), Jatani, India 33institutetext: Astrophysics Reseach Center (ARCO), The Open University of Israel, Ra’anana, 43107, Israel 44institutetext: Frontier Research Institute for Interdisciplinary Sciences, Tohoku University, 6-3 Aramaki aza Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 55institutetext: Department of Geophysics, Graduate School of Science, Tohoku University, 6-3 Aramaki aza Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 66institutetext: Institute of Space and Astronautical Science, Japan Aerospace Exploration Agency, 3-1-1 Yoshinodai, Chuo-ku, Sagamihara, Kanagawa 252-5210, Japan 77institutetext: Cluster for Pioneering Research, RIKEN, 2-1 Hirosawa, Wako, Saitama 351-0198, Japan 88institutetext: Department of Astronomy & Astrophysics, University of California, Santa Cruz, CA 95064, USA. 99institutetext: Department of Astronomy and Astrophysics, University of Chicago, Chicago, IL, USA 1010institutetext: School of Earth and Space Exploration, Arizona State University, Tempe, AZ 85287, USA 1111institutetext: Department of Physics, University of Warwick, Coventry CV4 7AL, UK 1212institutetext: Centre for Exoplanets and Habitability, University of Warwick, Coventry CV4 7AL, UK. 1313institutetext: Lowell Observatory, 1400 W Mars Hill Road, Flagstaff, AZ, 86001, USA 1414institutetext: Center for Computational Astrophysics, Flatiron Institute, New York, NY 10010, USA. 1515institutetext: Astrobiology Center, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan. 1616institutetext: National Astronomical Observatory of Japan,2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan. 1717institutetext: Department of Astronomy, Graduate University for Advanced Studies (SOKENDAI), Tokyo, 2-21-1 Osawa, Mitaka, Japan.
First Comparative Exoplanetology Within a Transiting Multi-planet System: Comparing the atmospheres of V1298 Tau b and c
The V1298 Tau system (20-30Myr), is a benchmark young multi-planet system that provides the opportunity to perform comparative exoplanetology between planets orbiting the same star right after their formation. We present the first atmospheric comparison between two planets in the same transiting system: V1298 Tau b and V1298 Tau c. We derive constraints on the mass of planet b and c (¡20M⊕ at 3 confidence level and respectively) and atmospheric metallicity (logZ/Z⊙=-2.04, -0.16 respectively) from atmospheric retrievals. The V1298 Tau planets, are likely to be similar in terms of mass at the current age, implying that both planets are potential sub-Neptune/super-Earth progenitors. However, planet c is expected to lose a higher fraction of its mass compared to planet b given its close proximity to the host star. Alternatively, the observed spectrum of planet c can be explained by atmospheric hazes, which is in contrast to planet b where efficient haze formation can be ruled out. Higher haze formation efficiency in planet c could be due to differences in atmospheric composition, temperature and higher UV flux incident compared to planet b.
1 Introduction
Demographic studies on mature exoplanets (Fulton et al., 2017; Van Eylen et al., 2018; Szabó & Kiss, 2011) suggest that early evolutionary processes like photoevaporation (Owen & Wu, 2017) and core-powered mass loss (Ginzburg et al., 2018; Gupta & Schlichting, 2019) significantly influence the structure and composition of super-Earths and sub-Neptunes. Therefore, observing planets in their early stages is key to understanding these processes.
V1298 Tau, a 20-30 Myr old weak-lined T-Tauri star in the foreground of the Taurus-Auriga star forming region (Oh et al., 2017; Luhman, 2018), is particularly valuable for such studies. It hosts four transiting planets, making it one of the youngest known multi-planet systems. The inner three planets (V1298 Tau c, d, and b) are in a near 3:2:1 mean-motion resonance (Feinstein et al., 2022; Sikora et al., 2023; David et al., 2019b), suggesting possible migration within a gas disk (Terquem & Papaloizou, 2018; Kajtazi et al., 2023). The orbital period of the outermost planet, V1298 Tau e, remains unconfirmed but is under investigation (Damasso et al., 2023).
In this paper, we compare the atmospheric composition and chemistry of two planets in the V1298 Tau system: V1298 Tau b (0.9 RJ, 24.14 day period) and V1298 Tau c (0.5 RJ, 8.24 day period). A comparison between the known physical properties of both planets is shown in Table 1. V1298 Tau c receives four times the stellar irradiation as V1298 Tau b due to their orbital distance ratio of 1:2. Stellar XUV flux is believed to play an important role in atmospheric chemistry (Kawashima & Ikoma, 2018; Tsai et al., 2023; Venot & Agúndez, 2015; Shulyak et al., 2020) as well as early evolution (Owen & Wu, 2013, 2017), which we aim to test with two planets in the same young system.
Mass measurement for these young planets is challenging due to stellar activity affecting radial velocity (RV) observations (Blunt et al., 2023). Initial RV estimates (Suárez Mascareño et al., 2021), which reported Jovian masses for the two outer V1298 Tau planets, has been challenged (Sikora et al., 2023; Finociety et al., 2023). Barat et al. (2024) estimated V1298 Tau b’s mass to be less than 23M⊕ at 3 level of confidence from HST/WFC3 transmission spectra, and ruled out earlier RV measurements at 5 level of confidence. Ongoing Transit timing variations (TTVs) studies (private communication) suggest that both V1298 Tau b and c are in the sub-Neptune/super-Earth mass regime, indicating they are likely sub-Neptune/super-Earth progenitors with similar masses.
We have observed a transmission spectrum of V1298 Tau c using the same instrumental configuration as Barat et al. (2024) to allow for comparison. We discuss the data reduction, light curve analysis and atmospheric modelling in Section 2, followed by the main results in Section 3. We discuss and interpret our findings in Section 4 and 5.
2 Data analysis and modelling
2.1 Observations
We observed one primary transit of V1298 Tau c (18th October 2021) with 8 HST orbits using the WFC3 instrument for Program GO 16462. The observations were taken with the G141 grism in bidirectional spatial scanning mode covering a range of 1.1-1.7 m. We used the pixel subarray and SPARS25, NSAMP=5 readout mode which resulted in 88.4 s exposures. We used a 0.23”/second scan rate resulting a 170 pixel long image in the spatial direction.
2.2 Data reduction and light curve analysis
We use a custom data reduction pipeline for analyzing the raw HST spatially scanned images (Arcangeli et al., 2018; Jacobs et al., 2022; Barat et al., 2024) mirroring the approach for planet b in Barat et al. (2024).
The broadband integrated light curve is shown in Figure 4. It shows ‘hook’ like systematics, which are typical for similar HST observations (Wakeford et al., 2016). Our observations covered only one pre-transit orbit and no ingress/egress observations. Typically, the first orbit of a visit for HST is known to exhibit large systematics, and it is common practice to not include it in time series analysis (e.g see Kreidberg et al., 2014; Jacobs et al., 2022; Barat et al., 2024). However, since we have only one pre-transit orbit, we do not remove it in our light curve analysis. Further details of the light curve fitting methods are provided in Appendix B.
2.3 Atmospheric modelling with ATMO and planet mass
We interpret the transmission spectra of V1298 Tau c and V1298 Tau b using a chemically consistent ATMO retrieval framework (Tremblin et al., 2015; Drummond et al., 2016; Goyal et al., 2018). ATMO has been previously used to analyze various exoplanet atmospheres (see for e.g Evans et al., 2018; Goyal et al., 2019; Carter et al., 2020). Details about the ATMO models are provided in Appendix D.
Figure 1 shows the best-fit atmospheric models for both planets alongside the observed spectra. The retrieved parameters are listed in Table 2, with posterior distributions in Figures 12, 13. Following Barat et al. (2024), we include planet mass as a free parameter. For planet b, the posterior distribution peaks around 8M⊕ with a 3 upper limit of 20M, consistent with Barat et al. (2024). For planet c, the posterior peaks around 17M⊕ with a 3 upper limit of 50M⊕, consistent with previous RV upper limits (Finociety et al., 2023; Sikora et al., 2023). These results suggest both planets are likely in the sub-Neptune/super-Earth mass range.
Our retrievals indicate low atmospheric metallicity for both planets, with planet b having 0.1-0.01 solar and planet c consistent with solar. The temperature of planet b peaks around 850-900K (See Appendix E for further discussion on planet b’s temperature), while planet c’s temperature is unconstrained. No constraints were found on the C/O ratio or clouds for either planet.
We also performed retrievals with fixed masses to assess their impact on other parameters. For V1298 Tau c, we used 10 and 50M, and for planet b, we used 10 and 23M. The posterior distributions for these fixed mass retrievals are shown in Figures 15, 14, with retrieved parameters in Table 2. The overall conclusion of sub-solar/solar atmospheric metallicity for both planets remains valid even with fixed masses.
3 Results
3.1 The transmission spectrum of V1298 Tau c
We compared the observed transmission spectrum of V1298 Tau c to a flat line and found that models with atmospheric water absorption are slightly favored (2.5). To confirm the presence of water, we conducted a free atmospheric retrieval using the open-source code PetitRadtrans (Mollière et al., 2019). We fixed the planet mass and temperature to the median of the ATMO retrieval posteriors (Table 2) and included a grey cloud deck opacity. The posterior distribution from this fit is shown in Figure 11. While we did not find constraints on CH4, our models preferred the presence of water, with a lower limit for the water volume mixing ratio () at 3 confidence level.
Comparing the spectrum of V1298 Tau c to its sibling planet, V1298 Tau b, revealed that the transit depth uncertainties for V1298 Tau c are larger by a factor of about 1.5. This higher uncertainty is likely due to including the first orbit of the visit for V1298 Tau c, known to have larger wavelength-dependent systematics. Consequently, we did not apply the common-mode correction to the spectroscopic light curves of planet c, as we did for planet b. The common-mode correction could correct for the curvature of the time-series baseline, which then allowed us to fit a linear baseline to the spectroscopic light curves of planet b.
3.2 Relative transmission spectroscopy between planet b and c
In Figure 1 (lower panel), we present the first relative transmission spectrum between two planets within the same system. This relative spectrum is independent of the star’s radius and limb darkening. Stellar inhomogeneities are known to contaminate transmission spectra (Désert et al., 2011; Rackham et al., 2019), which has also been highlighted in recent JWST observations(May et al., 2023; Moran et al., 2023; Lim et al., 2023). A relative transmission spectrum between planets theoretically bypasses stellar contamination issues, assuming consistent stellar contamination across both transits. This enables a better relative comparison of abundances between the planets without precise information on the stellar properties, including metallicity, elemental abundance ratios, and limb-darkening. However, in the case of V1298 Tau, an active star, changes in the stellar surface between observation epochs may still introduce contamination to the relative spectrum extracted at different times.
| Planet | Parameter | Free mass retrieval (ATMO) | Fixed mass (10M⊕) | Fixed mass (50M⊕) |
| logZ [solar] | -0.16 | -0.02 | -0.001 | |
| Rp [] | 0.48 | 0.48 | 0.48 | |
| V1298 Tau c | Teq [K] | 947 | 921 | 985 |
| log | 0.26 | 0.73 | -0.11 | |
| C/O | 0.06 | 0.07 | 0.06 | |
| Mass [M⊕] | 17 | 10 (fixed) | 50 (fixed) | |
| Planet | Parameter | Free mass retrieval (ATMO) | Fixed mass (10M⊕) | Fixed mass (23M⊕) |
| logZ [solar] | -2.04 | -1.80 | -0.86 | |
| Rp [] | 0.84 | 0.84 | 0.84 | |
| V1298 Tau b | Teq [K] | 889 | 806 | 824 |
| log | 1.03 | 0.88 | 0.25 | |
| C/O | 0.08 | 0.05 | 0.04 | |
| Mass [M⊕] | 7.58 | 10 (fixed) | 23 (fixed) |
3.3 Haze formation in the atmosphere of V1298 Tau planets
Hazes have been found to be ubiquitous in exoplanet atmospheres (Kempton et al., 2023; Libby-Roberts et al., 2020a; Knutson et al., 2014). The temperature of V1298 Tau b and c is expected to be between 600-1000K and the host star is young and active, producing copious amounts of XUV radiation, which are important ingredients for haze formation (Yu et al., 2021).
We simulated the transmission spectrum of V1298 Tau b and c, including hazes.We assumed tholin like composition for the hazes (Khare et al., 1984). The simulations were done following the approach in Kawashima & Ikoma (2019). The details of the haze microphysics model are provided in Kawashima & Ikoma (2018). We assumed 20M⊕ and 23M⊕ (Barat et al., 2024) for planet c and b respectively. For further details see Appendix F. The simulated transmission spectrum for V1298 Tau b and c, assuming a 1% haze formation efficiency is shown in Figure 1. The effect of hazes in the HST bandpass is to damp the water absorption feature. Given the large observable water feature in V1298 Tau b’s spectrum, efficient haze formation can be ruled out for this planet. However, atmospheric models with hazes can explain the observed transmission spectrum of planet c. If we assume a lower mass for planet c (10M⊕), which is consistent with current estimates from TTV analysis, we would require even higher haze formation rates to explain the observed spectrum of V1298 Tau c.
4 Discussion
4.1 Relative orbital distances and stellar density
In multi-planet systems, it is possible to fix one planet’s semi-major axis based on another’s measured semi-major axis, since they are related to their orbital periods ratio. However, due to known TTVs in this system, we refrain from fixing the semi-major axis in this analysis. The semi-major axis ratio from independent fits to the white light curves of V1298 Tau b and c is 2.080.1 and agrees with the expected value from their known orbital periods. Assuming circular orbits, the semi-major axis can be used to estimate stellar density (Sandford & Kipping, 2017). From the fits to our white light curves, we estimate the stellar density from the planet b and planet c to be 0.700.01gm/cm3 and 0.670.1gm/cm3. These values align with the published stellar density for V1298 Tau (0.680.1gm/cm3, David et al. (2019b)). indicating that it is inflation compared to similar-mass main sequence stars.
4.2 Limb darkening of V1298 Tau
Modeling stellar limb darkening is crucial for deriving transmission spectra. Prior studies often employed fixed limb darkening for main-sequence stars based on theoretical stellar structure models (e.g, see Deming et al., 2013; Sing et al., 2016). Barat et al. (2024) fitted a linear limb darkening coefficient for V1298 Tau, revealing significant differences compared to theoretical models for a main-sequence star with a similar spectral type (K1). Higher limb darkening compared to theoretical expectations for mature stars has previously been noted for -Pictoris (Landman et al., 2024). The fitted limb darkening coefficients and a theoretical linear limb darkening model is shown in Figure 9. This constitutes the first direct measurement of limb darkening on a weak-lined T-Tauri star.
In Section 4.1, we highlighted that V1298 Tau’s density is lower than that of similar-mass main-sequence stars. Furthermore, it has a relatively high magnetic field strength of the order of 100-300 Gauss (Finociety et al., 2023). High magnetic fields have been proposed as a potential cause for discrepancies between observed and theoretical limb darkening models (Kostogryz et al., 2024). Therefore, a combination of structural differences and higher magnetic fields compared to main-sequence stars, could explain the differences we find with the model.
While leveraging light curves from multiple planets in the same system could potentially reduce limb darkening coefficient uncertainties, the absence of ingress or egress coverage in the light curves of planet c led us to fix the limb darkening coefficients to the values from Barat et al. (2024). A comparison of the transmission spectrum of planet c derived using fixed and free limb darkening is shown in Figure 8. The transit depth uncertainties are higher by 5-10% for each spectroscopic bin in the latter case. Residuals in both cases are approximately 1.2 times the expected photon noise, indicating that stellar variability and instrumental systematics from the first orbit are likely to be the dominant sources of noise in our case.
4.3 Comparative exoplanetology between V1298 Tau b and c
A comparison between the atmospheres of V1298 Tau b and c, measured relative to atmospheric scale height (Figure 1, upper panel), shows that these two planets could host similar atmospheres. Our retrievals show that these planets are likely to have similar masses in the sub-Neptune/super-Earth regime (Table 2). Furthermore, we also find that the atmospheric metallicities of both the planets are solar/sub-solar, which is relatively low compared to their mature counterparts which have been found to have metal-rich/hazy (¿solar) atmospheres (Désert et al., 2011; Kempton et al., 2023; Madhusudhan et al., 2023; Libby-Roberts et al., 2020b). In this scenario, both the V1298 Tau planets are similar in terms of mass and composition, but their observable atmospheres are distinct from their mature counterparts.
Atmospheric hazes can be ruled out for V1298 Tau b, given the large atmospheric absorption feature, however for planet c atmospheric hazes can explain the observed spectrum.. A comparison between the transmission spectra of planet b and c shows that planet c could have a higher haze formation. Haze formation has been linked to the UV flux of the host star; a higher UV flux results in higher monomer production rates for organic hazes (Kawashima & Ikoma, 2019). Higher atmospheric metallicity could also lead to higher haze formation. The higher haze formation in planet c compared to its sibling planet could be due to difference in metallicity, temperature and incident UV flux of the two planets.
4.4 Comparing the early evolution of V1298 Tau b and c
We ran evolutionary models for the V1298 Tau planets. These models start from the time of disk dispersal, and model the structural and thermal evolution of the planet, including atmospheric mass loss due to photoevaporation (Owen & Wu, 2013). We run models assuming both a core-envelope interior structure, as well as with a gradually mixed interior. We have used the evolutionary models presented in Vazan et al. (2022). A brief description of the models is provided in Appendix G. Figure 2 shows a possible mass and radius evolution tracks for these planets.
V1298 Tau c being closer to its host star compared to planet b, is more susceptible to atmospheric mass loss driven by photoevaporation. While the final mass and radius depend on photoevaporation efficiency and the star’s activity level (Poppenhaeger et al., 2021), our models show that V1298 Tau c is likely to lose a larger fraction of its current envelope compared to planet b. Therefore, even if both planets have similar masses now, their early evolution could lead them to different mass-radius regimes.
Our models assume solar metallicity for the envelope opacity. However, the transmission spectrum of V1298 Tau c could also be explained using atmospheric hazes, unlike V1298 Tau b (Figure 1. Young planets are expected to cool and contract over time (Kubyshkina et al., 2020; Linder et al., 2019). The presence of high altitude hazes can increasing the atmospheric opacity, therefore increasing the cooling and contraction timescale (Lee & Chiang, 2015), leading to higher photoevaporative mass loss. Thus, a hazy atmosphere of V1298 Tau c as opposed to a clear atmosphere of V1298 Tau b, could lead to slower cooling and higher mass loss.
4.5 Comparing the composition of V1298 Tau planet with the exoplanet population
Figure 3 shows V1298 Tau b and c on the mass-metallicity and metallicity-orbital distance planes. The observed low atmospheric metallicity for these planets deviates from the mass-metallicity trend predicted by core-accretion models (Pollack et al., 1996; Bodenheimer & Pollack, 1986) and reported for exoplanets (Welbanks et al., 2019; Thorngren & Fortney, 2019). This could potentially be explained by ongoing evolutionary processes that transform their current metal-poor atmospheres into the metal-rich atmospheres seen in mature sub-Neptune/super-Earth planets (Désert et al., 2011; Kempton et al., 2023; Libby-Roberts et al., 2020a; Madhusudhan et al., 2023).
We compare the V1298 Tau system with the 30 Myr old HR 8799 system (Marois et al., 2010) with four directly imaged planets (HR8799 b, c , d and e) with masses 7, 10, 10, 10 (Marois et al., 2010; GRAVITY Collaboration et al., 2019). Nasedkin et al. (2024) using ground based spectra from the VLTI/GRAVITY program, report super-stellar(100 times stellar) atmospheric metallicity for these planets. Assuming a -0.65 dex metallicity for HR8799 (Swastik et al., 2021) we convert the retrieved atmospheric metallicity to stellar unit.
The HR 8799 planets exhibit higher atmospheric metallicity compared to the mass-metallicity trend (Thorngren & Fortney, 2019) predicted by core-accretion formation, while V1298 Tau planets fall below this trend. HR 8799 is a system of widely separated massive gas giants, whereas V1298 Tau likely consists of compact, low-mass sub-Neptune/super-Earth progenitors. Both systems likely host primordial atmospheres. The different mass regimes suggest HR 8799 planets underwent runaway gas accretion or gravitational collapse, unlike V1298 Tau planets. Theories propose close-in sub-Neptunes and super-Earths formed in situ within the water ice-line in dusty or depleting disks to prevent runaway gas accretion (Lee et al., 2014; Lee & Chiang, 2016). Wide orbit gas giants likely formed beyond the ice lines, allowing for runaway accretion (Pollack et al., 1996) or gravitational instability (Boss, 1997; Dodson-Robinson et al., 2009), but these mechanisms alone do not explain HR 8799’s high metallicity, necessitating the accretion of volatile-rich solids (Nasedkin et al., 2024). Formation location relative to ice lines could lead to differences in volatile enrichment, as beyond the ice-line, the disk contains volatile-rich solids, enhancing the metal content of primordial atmospheres (Öberg et al., 2011; Bitsch et al., 2019).
4.6 Formation and evolution of multi-planet systems
The mass of the V1298 Tau planets indicates they are likely sub-Neptune/super-Earth progenitors, which are in a near mean-motion resonance (1:2:3). Atmospheric retrievals suggest these planets could have low metallicity atmospheres. Population synthesis studies (Mordasini, 2018; Benz et al., 2014) show sub-Neptunes and super-Earths often form in multi-planet systems in mean motion resonance (Lambrechts et al., 2019). Emsenhuber et al. (2023) classified systems with multiple low-mass planets into two types: Class I: inner dry/rocky planets formed in situ with outer icy sub-Neptunes migrating inwards, and Class II: multiple sub-Neptunes forming beyond the water ice line and migrating inwards. In the first scenario, inner planets have volatile-poor atmospheres, whereas, in the second, all planets are volatile-rich.
The relatively low atmospheric metallicity of V1298 Tau b and c suggests a formation similar to class I planets . However, these young planets’ atmospheric composition may evolve. In the pebble accretion scenario, a metallicity gradient forms in the protoplanet’s envelope (Ormel et al., 2021; Bodenheimer et al., 2018), with metal-rich layers hidden deep inside. Over time, atmospheric mass loss can strip volatile-poor upper layers, enriching atmospheric metallicity (Vazan et al., 2022), aligning them with mature sub-Neptune/super-Earth population. In this case, the V1298 Tau system could also have a formation similar to the class II systems.
If V1298 Tau c has a hazy atmosphere, it might be more metal-rich than the outer planet b, which doesn’t fit either class I or class II systems. Class I systems have inner volatile-poor planets, while class II systems have planets with similar compositions. Due to its proximity to the star, V1298 Tau c is likely to lose more of its primordial envelope compared to V1298 Tau b 2. Thus, V1298 Tau b and c might have formed like class II systems, with volatiles hidden deep inside, but atmospheric mass loss caused planet c’s composition to evolve faster than planet b, explaining their difference in their atmospheric composition.
5 Conclusion
In this paper, we perform the first comparative study of atmospheres of multiple planets in the same transiting system. Comparing the transmission spectra of the young V1298 Tau b and c planets reveals that planet b has a large, haze-free envelope, while the presence of hazes in planet c cannot be ruled out. Atmospheric models suggest similar masses for both planets in the sub-Neptune/super-Earth regime. These models also find relatively low atmospheric metallicities for both these young planets compared to their matured counterparts, indicating they are sub-Neptune/super-Earth progenitors. Alternatively, if planet c has a hazy atmosphere, higher haze formation in planet c compared to planet b could be due to differences in composition, temperature, and incoming stellar UV flux. The higher envelope opacity of planet c would affect its cooling rate and evolution differently than for planet b.
6 Acknowledgements
J.M.D acknowledges support from the Amsterdam Academic Alliance (AAA) Program, and the European Research Council (ERC) European Union’s Horizon 2020 research and innovation program (grant agreement no. 679633; Exo-Atmos). This work is part of the research program VIDI New Frontiers in Exoplanetary Climatology with project number 614.001.601, which is (partly) financed by the Dutch Research Council (NWO). Y.K. acknowledges the support from JSPS KAKENHI Grant Numbers JP21K13984 and 22H05150.
7 Softwares used
numpy, scipy, matplotlib, emcee, lmfit, pandas, ATMO, PetitRadtrans
Appendix A Description of raw data reduction
We start our analysis from the ‘ima.fits’ files downloaded from MAST archive. During each exposure, the detector is read multiple times, without flushing out the accumulated charge (non-destructive reads). We create sub-exposures by subtracting consecutive non-destructive reads for each exposure which are reduced separately. This technique has been widely applied in past observations with HST (e.g, see Kreidberg et al. 2014; Arcangeli et al. 2018; Jacobs et al. 2022). We apply a wavelength dependent flat-field correction and flag bad pixels with data quality DQ=4, 32, or 512 by calwf3. We apply a local median filter to identify cosmic rays and clip pixels which are more than five median deviations from the median. To account for the dispersion direction drift of the spectrum (Wakeford et al. 2016) we use the first exposure of a visit as a template and shift the spectrum for an exposure along the dispersion direction to match the template. The maximum shift that we require is 0.1 pixels. We apply optimal extraction algorithm (Horne 1986) on each subexposure to maximize signal-to-noise ratio.
Appendix B Light curve fitting
HST WFC3 light curves often exhibit strong systematics like charge-trapping ramps and visit-long slopes (Kreidberg et al. 2014; Deming et al. 2013; Tsiaras et al. 2016; Wakeford et al. 2016). Common practice involves discarding the first orbit due to significantly larger systematics (See Figure 4) compared to the rest of the visit (e.g see, Wakeford et al. 2013; Arcangeli et al. 2018; Jacobs et al. 2022; Barat et al. 2024). However, in this case, we could not remove from the analysis because the transit occurred two hours earlier than expected from the linear ephemeris (Feinstein et al. 2022) due to large TTVs and we could only secure one orbit prior to ingress.
The first-orbit systematics are strongly wavelength dependent (Zhou et al. 2020, 2017) and cannot be corrected for using common-mode approaches, which had been previously applied to similar observations of V1298 tau b (Barat et al. 2024). Therefore, to handle these systematics, we used the physically motivated RECTE charge-trapping model (Zhou et al. 2017), which fits the first orbit ramp better than other methods (Berta et al. 2012; Kreidberg et al. 2014). We fixed the mid-transit time using simultaneous TESS observations (Feinstein et al. 2022). V1298 Tau shows strong rotational variability, with an average amplitude of 2 and a rotational period of 2.83 days (David et al. 2019b). We modeled it with a second-order polynomial for the out-of-transit baseline. See Appendix C for more discussion on the baseline modelling. The planetary transit signal was modelled using batman (Kreidberg 2015).
For the white light curve (1.12-1.6 m), we fit the RECTE parameters and the baseline polynomial coefficients, along with the planet’s transit depth and semi-major axis. Other orbital parameters were fixed to literature values (Feinstein et al. 2022; Sikora et al. 2023; David et al. 2019a). The RECTE parameters constitute 4 parameters: the number of initially filled fast/slow charge traps and the number of fast/slow charge traps that are filled in between orbits. Due to the lack of ingress and egress data, our limb darkening precision was impaired. However, by using the limb darkening parameter retrieved from the observations of planet b (Barat et al. 2024), we were able to significantly reduce the uncertainties on the transit spectrum (Figure 8). This highlights the advantage and uniqueness of studying multiple planets in the same system. For a discussion on the limb darkening of V1298 Tau see Appendix 4.2. We also independently measure the stellar density from the HST white light curves which is discussed in Appendix 4.1.
The fit was performed using emcee (Foreman-Mackey et al. 2013). The posterior distribution from these fits is shown in Supplementary Figure 5 and the best-fit transit model and residuals, are shown in Supplementary Figure 4 (upper panel).
Spectroscopic light curves were analyzed similarly by dividing the stellar spectra into 7-pixel bins, resulting in 17 channels (identical to V1298 Tau b). We fixed the semi-major axis from the white light curve fits and fit the RECTE parameters and baseline for each bin, and the transit depth. We fix the wavelength dependent limb-darkening coefficients to values derived in Barat et al. (2024). The detrended spectroscopic light curves and residuals are shown in Supplementary Figure 6.
Appendix C Modelling the baseline
Young stars exhibit quasi-periodic flux modulation due to rotational variability (e.g see Feinstein et al. 2022; David et al. 2019b; Mann et al. 2016, 2022). The transit of planet c analyzed in this work occurred near an inflection point of the stellar baseline, resulting in significant deviation from a linear function. We modeled the baseline using linear, quadratic, and cubic polynomial functions for all the spectroscopic channels. The transit depths from all three fits were consistent within 1, but quadratic and cubic polynomials were statistically favored over a linear baseline (See Supplementary Figure 7).
Higher-order polynomials, however, introduce correlation between the transit depth and baseline, increasing uncertainties and scatter in the transmission spectrum. The Bayesian Information Criterion (BIC) did not improve significantly with a cubic baseline compared to a quadratic one, and the cubic polynomial coefficients strongly correlated with the transit depths, likely adding noise. Therefore, we used a quadratic polynomial to model the baseline for the final transmission spectrum.
Appendix D Atmospheric models with ATMO
ATMO’s python based retrieval framework was applied and bench-marked in Lewis et al. (2020) and Rathcke et al. (2021) for HAT-P-41b and WASP-79b, respectively . This same retrieval framework is applied to V1298 Tau b and c in this work. This framework includes a Nested sampling Bayesian sampler DYNESTY (Speagle 2020) coupled to ATMO. We perform chemical equilibrium retrieval with isothermal - profile. Chemical equilibrium abundances are computed using Gibbs energy minimization as detailed in Drummond et al. (2016). We include H2-H2, H2-He Collision Induced opacity (CIA), as well as H2O, CH4 and CO high temperature molecular opacity in the retrieval. The clouds are treated as grey cloud deck between 1 and 50 millibar pressure levels, but with variable cloud opacity governed by the cloud factor used as the variable in the retrieval, as detailed in Goyal et al. (2018).
Appendix E Comparing V1298 Tau b ATMO retrieval with Barat et al. (2024)
For this study, we performed atmospheric retrieval on V1298 Tau b using the same setup as for planet c, assuming an isothermal atmosphere. Our results, indicating sub-solar metallicity and a mass upper limit of 20M⊕, align with Barat et al. (2024), who used PetitRadtrans. However, our retrieved temperatures are higher (800-900K) compared to their 600-700K. The discrepancy arises because Barat et al. (2024) used a non-isothermal temperature-pressure (T-P) profile with high interior temperature and vertical mixing to explain the non-detection of methane for this planet. Our isothermal model lacks this flexibility, leading to higher temperatures beyond the CH4/CO transition (Fortney et al. 2020) to eliminate methane from the upper atmosphere.
Appendix F Haze formation models
The haze microphysics models were simulated using the models described in Kawashima & Ikoma (2018), following the approach of Kawashima & Ikoma (2019). First, using a photochemical model the abundance of gaseous species was calculated. The XUV spectrum of V1298 Tau was adopted from Duvvuri et al. (2023). Haze production rates (or efficiency of haze formation) are uncertain, therefore we assumed a certain fraction of precursor molecules to be dissociated to form haze monomers. The haze precursor molecules are assumed to be CH4, HCN and C2H2 in our model. We derived vertical density and radius distributions of the haze particles following Kawashima & Ikoma (2018). We assumed solar elemental abundance ratio and other parameters such as eddy diffusion coefficient, monomer radius, internal density, and refractive index of haze particles similar to Kreidberg et al. (2022). We simulated transmission spectra, for V1298 Tau b and c assuming 23 and 20M⊕ respectively with obtained profiles of gas and haze particles. In Figure 1 we show the simulated transmission spectra assuming a 1 haze production efficiency.
Appendix G Theoretical evolutionary models
Our models start from the time of disk dispersal and calculate the thermal evolution and structural evolution. We consider both core-envelope as well as gradually mixed interior structures. We include atmospheric mas loss due to photoevaporation from the topmost layer of the planets using the formalism of Rogers & Owen (2021). The mass loss rates are estimated from an energy-limited model (Erkaev et al. 2007), assuming a fixed photoevaporation efficiency. Our models include core erosion and mixing with the envelope. The evolutionary models are described in Vazan et al. (2022). We choose the initial conditions (mass, radius) such that it matches the current mass and radius for the planets. We assumed 23M⊕ and 20M⊕ for the current mass estimates of the V1298 Tau b and c respectively. We run simulations with both core-envelope and gradually mixed interior models. Both interior structure models can explain the current mass/radius of the planets. Mass and radius (normalized at current value) evolution models for the V1298 Tau planets are shown in Figure 2.
Appendix H Effect of stellar activity on transmission spectrum
We estimate the effect of stellar activity on the transmission spectrum of V1298 Tau c. We adopt the method outlined in Rackham et al. (2019), where we estimate a correction factor based on the spot coverage fraction and spot temperature contrast. We adopt a spot coverage fraction of 20% for V1298 Tau (Feinstein et al. 2021). We adopt an extreme spot contrast of 1000K following spot contrasts derived for a sample of T Tauri stars (Koen 2016). We show a comparison between the observed and corrected transmission spectrum in Supplementary Figure 10. The observed and contamination corrected normalized spectra are consistent within 1 for all spectroscopic channels.
Since V1298 Tau is a K-type star, its contamination function does not show molecular features like M-dwarfs (Barclay et al. 2021). Therefore, the effect of the stellar correction is to change the level of the continuum in this case. However, the atmospheric features, which are relative to the continuum remain largely unaffected.
Appendix I Supplementary Figures
| Planet name | Mass [M⊕] | Atmospheric metallicity [solar] | Semi-major axis [AU] | Stellar metallicity [dex] | Reference |
| Wasp-18b | 10.200.35 | 1.03 | 0.02 | 0.110.08 | Coulombe et al. (2023) |
| Wasp-39b | 0.2800.03 | 10 | 0.048 | -0.030.1 | Rustamkulov et al. (2023) |
| Wasp-77Ab | 1.660.06 | 0.12 | 0.023 | -0.010.1 | August et al. (2023) |
| Wasp-96b | 0.480.03 | 7.5 | 0.045 | 0.140.19 | Radica et al. (2023) |
| TOI-270d | 0.0150.001 | 225 | 0.07 | -0.20.1 | Benneke et al. (2024) |
| Wasp-107b | 0.10.005 | 10 | 0.05 | 0.020.09 | Welbanks et al. (2024) |
| HD149026b | 0.3580.01 | 20 | 0.0436 | 0.330.1 | Gagnebin et al. (2024) |
| HD209458b | 0.59029040.01 | 3 | 0.05 | -0.010.05 | Xue et al. (2024) |
| Wasp-80b | 0.530.03 | 6 | 0.03 | -0.10.1 | Bell et al. (2023) |
| Wasp-17b | 0.780.2 | 50 | 0.051 | -0.070.1 | Grant et al. (2023) |
| GJ1214b | 0.0260.001 | 100 | 0.01 | 0.020.1 | Kempton et al. (2023) |

### v2

@@
 First Comparative Exoplanetology Within a Transiting Multi-planet System: Comparing the atmospheres of V1298 Tau b and c
-11email: s.barat@uva.nl 22institutetext: School of Earth and Planetary Sciences (SEPS), National Institute of Science Education and Research (NISER), Jatani, India 33institutetext: Astrophysics Reseach Center (ARCO), The Open University of Israel, Ra’anana, 43107, Israel 44institutetext: Frontier Research Institute for Interdisciplinary Sciences, Tohoku University, 6-3 Aramaki aza Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 55institutetext: Department of Geophysics, Graduate School of Science, Tohoku University, 6-3 Aramaki aza Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 66institutetext: Institute of Space and Astronautical Science, Japan Aerospace Exploration Agency, 3-1-1 Yoshinodai, Chuo-ku, Sagamihara, Kanagawa 252-5210, Japan 77institutetext: Cluster for Pioneering Research, RIKEN, 2-1 Hirosawa, Wako, Saitama 351-0198, Japan 88institutetext: Department of Astronomy & Astrophysics, University of California, Santa Cruz, CA 95064, USA.
+11email: s.barat@uva.nl 22institutetext: School of Earth and Planetary Sciences (SEPS), National Institute of Science Education and Research (NISER), Jatani, India 33institutetext: Astrophysics Research Center (ARCO), The Open University of Israel, Ra’anana, 43107, Israel 44institutetext: Frontier Research Institute for Interdisciplinary Sciences, Tohoku University, 6-3 Aramaki aza Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 55institutetext: Department of Geophysics, Graduate School of Science, Tohoku University, 6-3 Aramaki aza Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 66institutetext: Institute of Space and Astronautical Science, Japan Aerospace Exploration Agency, 3-1-1 Yoshinodai, Chuo-ku, Sagamihara, Kanagawa 252-5210, Japan 77institutetext: Cluster for Pioneering Research, RIKEN, 2-1 Hirosawa, Wako, Saitama 351-0198, Japan 88institutetext: Department of Astronomy & Astrophysics, University of California, Santa Cruz, CA 95064, USA.
 99institutetext: Department of Astronomy and Astrophysics, University of Chicago, Chicago, IL, USA 1010institutetext: School of Earth and Space Exploration, Arizona State University, Tempe, AZ 85287, USA 1111institutetext: Department of Physics, University of Warwick, Coventry CV4 7AL, UK 1212institutetext: Centre for Exoplanets and Habitability, University of Warwick, Coventry CV4 7AL, UK.
@@
 1616institutetext: National Astronomical Observatory of Japan, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan.
-1717institutetext: Department of Astronomy, Graduate University for Advanced Studies (SOKENDAI), Tokyo, 2-21-1 Osawa, Mitaka, Japan.
+1717institutetext: Department of Astronomical Science, The Graduate University for Advanced Studies, SOKENDAI, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan.
 First Comparative Exoplanetology Within a Transiting Multi-planet System: Comparing the atmospheres of V1298 Tau b and c
-The V1298 Tau system (20-30Myr), is a benchmark young multi-planet system that provides the opportunity to perform comparative exoplanetology between planets orbiting the same star right after their formation.
-We present the first atmospheric comparison between two planets in the same transiting system: V1298 Tau b and V1298 Tau c.
-We derive constraints on the mass of planet b and c (¡20M⊕ at 3 confidence level and respectively) and atmospheric metallicity (logZ/Z⊙=-2.04, -0.16 respectively) from atmospheric retrievals.
-The V1298 Tau planets, are likely to be similar in terms of mass at the current age, implying that both planets are potential sub-Neptune/super-Earth progenitors.
-However, planet c is expected to lose a higher fraction of its mass compared to planet b given its close proximity to the host star.
-Alternatively, the observed spectrum of planet c can be explained by atmospheric hazes, which is in contrast to planet b where efficient haze formation can be ruled out.
-Higher haze formation efficiency in planet c could be due to differences in atmospheric composition, temperature and higher UV flux incident compared to planet b.
+The V1298 Tau system is a multi-planet system that provides the opportunity to perform comparative exoplanetology between planets orbiting the same star.
+Because of its young age (20-30 Myr), this system also provides the opportunity to compare the planet’s early evolutionary properties, right after their formation.
+We present the first atmospheric comparison between two transiting exoplanets within the same multiple planet system: V1298 Tau b and V1298 Tau c.
+We observed one primary transit for each planet with the Hubble Space Telescope (HST), using Grism 141 (G141) of Wide Field Camera 3 (WFC3).
+We fit the spectroscopic light curves using state-of-the-art techniques to derive the transmission spectrum for planet c and adopt the transmission spectrum of planet b obtained with the same observing configuration and data analysis methods from previous studies.
+We measure the mass of planet b and c (, M⊕; respectively) from the transmission spectrum and find the two planets to have masses in the Neptune/sub-Neptune regime.
+Using atmospheric retrievals, we measure and compare the atmospheric metallicities of planet b and c (logZ/Z⊙=-2.04, logZ/Z⊙= -0.16, respectively), and find them consistent with solar/sub-solar, which is low (at least 1 order or magnitude) compared to known mature Neptune/sub-Neptune planets.
+This discrepancy could be explained by ongoing early evolutionary mechanisms, which are expected to enrich the atmospheres of such young planets as they mature.
+Alternatively, the observed spectrum of planet c can be explained by atmospheric hazes, which is in contrast to planet b, where efficient haze formation can be ruled out.
+Higher haze formation efficiency in planet c could be due to differences in atmospheric composition, temperature and higher UV flux compared to planet b.
+In addition, planet c is likely to experience a higher fraction of mass loss compared to planet b, given its proximity to the host star.
 1 Introduction
-Demographic studies on mature exoplanets (Fulton et al., 2017; Van Eylen et al., 2018; Szabó & Kiss, 2011) suggest that early evolutionary processes like photoevaporation (Owen & Wu, 2017) and core-powered mass loss (Ginzburg et al., 2018; Gupta & Schlichting, 2019) significantly influence the structure and composition of super-Earths and sub-Neptunes.
-Therefore, observing planets in their early stages is key to understanding these processes.
-V1298 Tau, a 20-30 Myr old weak-lined T-Tauri star in the foreground of the Taurus-Auriga star forming region (Oh et al., 2017; Luhman, 2018), is particularly valuable for such studies.
-It hosts four transiting planets, making it one of the youngest known multi-planet systems.
-The inner three planets (V1298 Tau c, d, and b) are in a near 3:2:1 mean-motion resonance (Feinstein et al., 2022; Sikora et al., 2023; David et al., 2019b), suggesting possible migration within a gas disk (Terquem & Papaloizou, 2018; Kajtazi et al., 2023).
-The orbital period of the outermost planet, V1298 Tau e, remains unconfirmed but is under investigation (Damasso et al., 2023).
-In this paper, we compare the atmospheric composition and chemistry of two planets in the V1298 Tau system: V1298 Tau b (0.9 RJ, 24.14 day period) and V1298 Tau c (0.5 RJ, 8.24 day period).
-A comparison between the known physical properties of both planets is shown in Table 1.
+Demographic studies on mature exoplanets (Szabó & Kiss 2011; Fulton et al.
+2017; Fulton & Petigura 2018; Van Eylen et al.
+2018) suggest that early evolutionary processes like photoevaporation (Owen & Wu 2017) and core-powered mass loss (Ginzburg et al.
+2018; Gupta & Schlichting 2019) significantly influence the structure and composition of low mass planets, such as Neptunes and sub-Neptunes.
+Therefore, observing planets right after their formation is the key to understanding these early evolutionary processes.
+V1298 Tau, a 20-30 Myr old weak-lined T-Tauri star in the foreground of the Taurus-Auriga star-forming region (Oh et al.
+2017; Luhman 2018), is particularly valuable for such studies.
+It has a temperature of 5000 K, with a radius and mass of 1.340.05 R⊙ and 1.10.05 M⊙, respectively, and has a spectral type between K0-K1 (David et al.
+2019a).
+It is a pre-main sequence star, and shows rotational variability at the level of 2% peak-to-peak and a rotational period of 2.860.01 days (David et al.
+2019a; Feinstein et al.
+2022; Sikora et al.
+2023).
+Such rotational variability is typical of young pre-main sequence stars (e.g.
+see David et al.
+2016; Mann et al.
+2016, 2022) and is due to the presence of spots and faculae on their photospheres.
+Feinstein et al.
+(2021) reproduced the optical photometric variability using a model with 20% spot coverage on the surface assuming a temperature contrast of 500 K between the photosphere and the spots.
+V1298 Tau hosts four transiting planets, making it one of the youngest known transiting multi-planet systems.
+The inner three planets (V1298 Tau c, d, and b) are in a near 3:2:1 mean-motion resonance (David et al.
+2019b; Feinstein et al.
+2022; Sikora et al.
+2023), suggesting possible migration within a gas disk (Terquem & Papaloizou 2018; Kajtazi et al.
+2023).
+The mutual inclinations between the planets in this system are known to be low (Johnson et al.
+2022), indicating planar orbital geometry for this system.
+The orbital period of the outermost planet, V1298 Tau e, remains unconfirmed but is under investigation (Damasso et al.
+2023).
+With a wealth of observations of exoplanet atmospheres, we have entered an era of understanding atmospheric processes, planet formation and early evolution through comparative studies.
+There have been previous studies aimed at understanding the physical and chemical processes of mature transiting exoplanet atmospheres from a statistical point of view.
+For example, the dependence of thermal structure and opacity sources in the atmosphere on insolation (e.g, see Garhart et al.
+2020; Baxter et al.
+2020), emergence of clouds in the atmosphere (Sing et al.
+2016; Fu et al.
+2017; Keating et al.
+2019), disequilibrium chemistry (Baxter et al.
+2021; Fortney et al.
+2020; Tsai et al.
+2023; Baeyens et al.
+2022) and the formation and composition of hazes (Gao et al.
+2020; Brande et al.
+2024) has been studied by comparing the atmospheres of different exoplanets.
+But, differences in the host stars can introduce biases due to uncertainty in their elemental abundances, activity levels and evolution history.
+These challenges can be mitigated by performing comparative exoplanetology between multiple planets in the same system, since they share their host star.
+In this paper, we compare the atmospheric metallicity, chemistry and potential evolution of two planets in the V1298 Tau system: V1298 Tau b (0.9 RJ, 24.14 day period) and V1298 Tau c (0.5 RJ, 8.24 day period) by studying their atmospheres.
+Given its young age, this system also allows for atmospheric comparison with mature transiting exoplanets as well as, widely separated directly imaged planets which have similar age.
+A comparison between the known physical properties of planet b and c is shown in Table 1.
 V1298 Tau c receives four times the stellar irradiation as V1298 Tau b due to their orbital distance ratio of 1:2.
-Stellar XUV flux is believed to play an important role in atmospheric chemistry (Kawashima & Ikoma, 2018; Tsai et al., 2023; Venot & Agúndez, 2015; Shulyak et al., 2020) as well as early evolution (Owen & Wu, 2013, 2017), which we aim to test with two planets in the same young system.
-Mass measurement for these young planets is challenging due to stellar activity affecting radial velocity (RV) observations (Blunt et al., 2023).
-Initial RV estimates (Suárez Mascareño et al., 2021), which reported Jovian masses for the two outer V1298 Tau planets, has been challenged (Sikora et al., 2023; Finociety et al., 2023).
+Stellar XUV flux is believed to play an important role in atmospheric chemistry (Venot & Agúndez 2015; Kawashima & Ikoma 2018; Shulyak et al.
+2020; Tsai et al.
+2023) as well as early evolution (Owen & Wu 2013, 2017), which we aim to test with two planets in the same young system.
+Mass measurement for these young planets is challenging due to stellar activity induced jitter affecting radial velocity (RV) observations (Brems et al.
+2019; Blunt et al.
+2023).
+Initial RV studies (Suárez Mascareño et al.
+2021) reported Jovian masses (20070, 400100 M⊕) for the two outer V1298 Tau planets (b and e, respectively) and mass upper limits (70 and 100 M⊕) for the inner two (c and d, respectively).
+Follow-up RV observations of this system have challenged these mass measurements (Sikora et al.
+2023; Finociety et al.
+2023).
+Finociety et al.
+(2023) reports an upper limit of 50 M⊕ for planet c and d and an upper limit of 150 M⊕ for planet b.
+Sikora et al.
+(2023) reports a mass of 2010 M⊕ for planet c and upper limits of 36 and 160 M⊕ for planets d and b, respectively.
+The follow-up RV papers are consistent with Suárez Mascareño et al.
+(2021) for the mass of V1298 Tau e, but do not agree with the mass of planet b from Suárez Mascareño et al.
+(2021).
 Barat et al.
-(2024) estimated V1298 Tau b’s mass to be less than 23M⊕ at 3 level of confidence from HST/WFC3 transmission spectra, and ruled out earlier RV measurements at 5 level of confidence.
-Ongoing Transit timing variations (TTVs) studies (private communication) suggest that both V1298 Tau b and c are in the sub-Neptune/super-Earth mass regime, indicating they are likely sub-Neptune/super-Earth progenitors with similar masses.
-We have observed a transmission spectrum of V1298 Tau c using the same instrumental configuration as Barat et al.
-(2024) to allow for comparison.
-We discuss the data reduction, light curve analysis and atmospheric modelling in Section 2, followed by the main results in Section 3.
+(2024) estimated V1298 Tau b’s mass to be less than 23 M⊕ at 3 level of confidence from HST/WFC3 transmission spectra, and ruled out RV estimate from Suárez Mascareño et al.
+(2021) at 5 level of confidence.
+Ongoing Transit timing variations (TTVs) studies (Livingston et al in prep) suggest that both V1298 Tau b and c are in the sub-Neptune mass regime.
+In this paper, we compare the transmission spectra of V1298 Tau b and c.
+For planet b, we adopt the spectrum obtained in Barat et al.
+(2024).
+We analyze the primary transit observations of V1298 Tau c, observed using the same instrumental configuration as Barat et al.
+(2024), and present the transmission spectrum of planet c.
+We discuss the data reduction and light curve analysis for the visit of V1298 Tau c in Section 2, followed by the main results in Section 3.
 We discuss and interpret our findings in Section 4 and 5.
-2 Data analysis and modelling
+| Parameter | V1298 Tau b | V1298 Tau c |
+| Period [days] 11111111Sikora et al. (2023) | 24.1404100.000022 | 8.2487200.000024 |
+| Semi-major axis [AU] | 0.17160.0028 | 0.08390.0014 |
+| Equilibrium temperature [K] | 68515 | 97921 |
+| Radius [R⊕] | 9.950.37 | 5.240.24 |
+| Known mass [M⊕] | 23 22222222Barat et al. (2024) | 2010 11111111Sikora et al. (2023) |
+2 Data analysis and modeling
 2.1 Observations
 We observed one primary transit of V1298 Tau c (18th October 2021) with 8 HST orbits using the WFC3 instrument for Program GO 16462.
-The observations were taken with the G141 grism in bidirectional spatial scanning mode covering a range of 1.1-1.7 m.
-We used the pixel subarray and SPARS25, NSAMP=5 readout mode which resulted in 88.4 s exposures.
-We used a 0.23”/second scan rate resulting a 170 pixel long image in the spatial direction.
+The observations were taken with the G141 in bidirectional spatial scanning mode covering a range of 1.1-1.7 m.
+We used the pixel subarray and SPARS25, NSAMP=5 readout mode, which resulted in 88.4 s exposures.
+We used a 0.23”/second scan rate, resulting in a 170 pixel long image in the spatial direction.
+The observations for V1298 Tau b had been taken using the same instrumental settings, with 10 HST orbits.
+See Barat et al.
+(2024) for further details.
 2.2 Data reduction and light curve analysis
-We use a custom data reduction pipeline for analyzing the raw HST spatially scanned images (Arcangeli et al., 2018; Jacobs et al., 2022; Barat et al., 2024) mirroring the approach for planet b in Barat et al.
+We use a custom data reduction pipeline for analyzing the raw HST spatially scanned images (Arcangeli et al.
+2018, 2019; Jacobs et al.
+2022; Barat et al.
+2024) for the V1298 Tau c visit.
+The same pipeline had been used for analyzing the visit of planet b in Barat et al.
 (2024).
-The broadband integrated light curve is shown in Figure 4.
-It shows ‘hook’ like systematics, which are typical for similar HST observations (Wakeford et al., 2016).
-Our observations covered only one pre-transit orbit and no ingress/egress observations.
-Typically, the first orbit of a visit for HST is known to exhibit large systematics, and it is common practice to not include it in time series analysis (e.g see Kreidberg et al., 2014; Jacobs et al., 2022; Barat et al., 2024).
-However, since we have only one pre-transit orbit, we do not remove it in our light curve analysis.
+We outline the data reduction methods we use in Appendix A.
+The broadband integrated white light curve for the visit of V1298 Tau c is shown in Figure 1.
+It shows ‘hook’ like systematics for each orbit, which have been observed for similar HST observations (Berta et al.
+2012; Deming et al.
+2013; Ranjan et al.
+2014).
+We also see that the orbital ramp is much larger in the first orbit compared to the rest of the orbits for this visit.
+This is a typical feature of HST visits where the first orbit shows larger systematics (Wakeford et al.
+2016; Arcangeli et al.
+2021).
+The light curves also show a long-term curvature, which is likely due to the rotational variability from active regions on the photosphere of this young star.
+This variability has been reported by Barat et al.
+(2024) and is well known for young pre-main sequence stars which have large spot coverage and fast rotation rates (e.g.
+see David et al.
+2016, 2019b; Mann et al.
+2022).
+Such variability is known to affect multi-epoch transmission spectra (e.g, Désert et al.
+2011b; Berta et al.
+2012; Rackham et al.
+2019).
+These observations of planet c were planned using linear ephemeris from Feinstein et al.
+(2022), but this system is known to exhibit TTVs and the actual transit occurred 2 hours prior to the predicted time from the linear ephemeris.
+Therefore, our observations covered only one pre-transit orbit and no exposures during ingress or egress.
+Typically, the first orbit of a visit for HST is known to exhibit large systematics, and it is common practice not to include it in time series analysis (e.g.
+see Berta et al.
+2012; Ranjan et al.
+2014; Jacobs et al.
+2022; Barat et al.
+2024).
+In the case of V1298 Tau b, there were three pre-transit orbits, and therefore, the first orbit could be removed from the analysis.
+However, since we have only one pre-transit orbit for the visit of V1298 Tau c, we could not remove the first orbit while analyzing its light curve.
+The white light curve of planet c has been modeled assuming a physically motivated charge-trapping model RECTE (Zhou et al.
+2017) to model the ‘hook-like’ orbital ramps and a second order polynomial function to model the baseline for the planet c visit.
+The RECTE model is known to fit the first orbit better compared to other known methods (e.g, Berta et al.
+2012).
+The planet transit was modeled using batman (Kreidberg 2015), with orbital parameters taken from Sikora et al.
+(2023).
+This same approach had been used to fit the white light curve of planet b (Barat et al.
+2024).
+A linear limb darkening model was assumed for V1298 Tau.
+Barat et al.
+(2024) fit for the linear limb darkening coefficient for the visit of planet b.
+Since the planet c visit does not have any ingress/egress coverage, we have used the fitted limb darkening for V1298 Tau from Barat et al.
+(2024) as a fixed parameter for the analysis of the planet c visit (Figure 10).
+V1298 Tau b and c have impact parameters (0.34 and 0.46, respectively) consistent within 1 of each other (David et al.
+2019b, a).
+Furthermore, Johnson et al.
+(2022) report a relatively low mutual inclination (019∘), indicating a planar orbital geometry of the V1298 Tau system.
+Therefore, both planets would transit a similar part of the stellar photosphere, and are expected to to show similar limb darkening during their primary transits.
 Further details of the light curve fitting methods are provided in Appendix B.
-2.3 Atmospheric modelling with ATMO and planet mass
-We interpret the transmission spectra of V1298 Tau c and V1298 Tau b using a chemically consistent ATMO retrieval framework (Tremblin et al., 2015; Drummond et al., 2016; Goyal et al., 2018).
-ATMO has been previously used to analyze various exoplanet atmospheres (see for e.g Evans et al., 2018; Goyal et al., 2019; Carter et al., 2020).
-Details about the ATMO models are provided in Appendix D.
-Figure 1 shows the best-fit atmospheric models for both planets alongside the observed spectra.
-The retrieved parameters are listed in Table 2, with posterior distributions in Figures 12, 13.
-Following Barat et al.
-(2024), we include planet mass as a free parameter.
-For planet b, the posterior distribution peaks around 8M⊕ with a 3 upper limit of 20M, consistent with Barat et al.
+We obtain 17 spectroscopic light curves for the planet c visit by using 7-pixel bins.
+The same binning had been used for planet b in Barat et al.
 (2024).
-For planet c, the posterior peaks around 17M⊕ with a 3 upper limit of 50M⊕, consistent with previous RV upper limits (Finociety et al., 2023; Sikora et al., 2023).
-These results suggest both planets are likely in the sub-Neptune/super-Earth mass range.
-Our retrievals indicate low atmospheric metallicity for both planets, with planet b having 0.1-0.01 solar and planet c consistent with solar.
-The temperature of planet b peaks around 850-900K (See Appendix E for further discussion on planet b’s temperature), while planet c’s temperature is unconstrained.
-No constraints were found on the C/O ratio or clouds for either planet.
-We also performed retrievals with fixed masses to assess their impact on other parameters.
-For V1298 Tau c, we used 10 and 50M, and for planet b, we used 10 and 23M.
-The posterior distributions for these fixed mass retrievals are shown in Figures 15, 14, with retrieved parameters in Table 2.
-The overall conclusion of sub-solar/solar atmospheric metallicity for both planets remains valid even with fixed masses.
+For the analysis of spectroscopic light curves of planet c, the semi-major axis is fixed from the white light curve fit of planet c.
+This approach had also been followed for the planet b visit.
+In the analysis of planet b (Barat et al.
+2024), the common-mode method (using white light curve systematics models to correct the spectroscopic light curves) was applied to de-trend the spectroscopic light curves, assuming that the systematics were achromatic.
+However, the first orbit is known to exhibit wavelength-dependent systematics (Zhou et al.
+2017, 2020), and therefore in the analysis of V1298 Tau c we could not use the common-mode technique for de-trending spectroscopic light curves.
+We used RECTE (Zhou et al.
+2017) to model the ‘hook-like’ orbital ramps and a second order polynomial function to model the baseline of the spectroscopic light curves for the planet c visit.
+The parameters for the RECTE as well as the polynomial coefficients, were fitted for each spectroscopic light curve.
+We tested different models for the baseline (linear, quadratic and cubic).
+We found that the quadratic model fits the observed baseline the best.
+See Appendix C for further details on the baseline model.
+The limb darkening coefficients were fixed from Barat et al.
+(2024).
+We derive the transmission spectrum of V1298 Tau c following the approach outlined in this Section, and we present it in Figure 2, along with the spectra of planet b, which has been adopted from Barat et al.
+(2024).
+2.3 Atmospheric models from ATMO
+We interpret the transmission spectrum of V1298 Tau c using the equilibrium chemistry ATMO retrieval framework (Tremblin et al.
+2015; Drummond et al.
+2016; Goyal et al.
+2018).
+In this paper, for a consistent comparison with planet c, we also reanalyze the spectrum of V1298 Tau b adopted from Barat et al.
+(2024), using the same framework as applied for planet c.
+ATMO has been previously used to analyze various exoplanet atmospheres (see for e.g Evans et al.
+2018; Goyal et al.
+2019; Carter et al.
+2020).
+ATMO’s python based retrieval framework was applied and benchmarked in Lewis et al.
+(2020) and Rathcke et al.
+(2021) for HAT-P-41b and WASP-79b, respectively.
+This same retrieval framework is applied to V1298 Tau b and c in this work.
+This framework includes a Nested sampling Bayesian sampler DYNESTY (Speagle 2020) coupled to ATMO.
+We model the planets using a 1D atmosphere with an isothermal T-P profile and calculate the molecular abundances, assuming chemical equilibrium.
+Atmospheric metallicity and C/O ratio are used as parameters to calculate the elemental abundances.
+For a given metallicity, in our model the C/O ratio is implemented by keeping the number of C atoms fixed, and adjusting the number of O atoms.
+We include a grey opacity source to model clouds.
+We include H2-H2, H2-He Collision Induced opacity (CIA), as well as H2O, CH4 and CO high temperature molecular opacity in the retrieval.
+These molecules are chosen, as the HST bandpass includes absorption features from these molecules.
+We include planet mass as a free parameter in our retrieval setup for both planets b and c.
+We put similar uniform priors on atmospheric metallicity (0.01 - 1000 solar), O/H ratio (0.01 - 100 solar), degree of cloudiness (0.1 - 100 Rayleigh scattering cross-section at 350 nm) for both V1298 Tau b and c.
+We put uniform prior on the mass between 5 - 50 M⊕ for planet c (Finociety et al.
+2023; Sikora et al.
+2023) and between 5-25 M⊕ for planet b (Barat et al.
+2024).
 3 Results
 3.1 The transmission spectrum of V1298 Tau c
-We compared the observed transmission spectrum of V1298 Tau c to a flat line and found that models with atmospheric water absorption are slightly favored (2.5).
-To confirm the presence of water, we conducted a free atmospheric retrieval using the open-source code PetitRadtrans (Mollière et al., 2019).
-We fixed the planet mass and temperature to the median of the ATMO retrieval posteriors (Table 2) and included a grey cloud deck opacity.
-The posterior distribution from this fit is shown in Figure 11.
-While we did not find constraints on CH4, our models preferred the presence of water, with a lower limit for the water volume mixing ratio () at 3 confidence level.
-Comparing the spectrum of V1298 Tau c to its sibling planet, V1298 Tau b, revealed that the transit depth uncertainties for V1298 Tau c are larger by a factor of about 1.5.
-This higher uncertainty is likely due to including the first orbit of the visit for V1298 Tau c, known to have larger wavelength-dependent systematics.
-Consequently, we did not apply the common-mode correction to the spectroscopic light curves of planet c, as we did for planet b.
-The common-mode correction could correct for the curvature of the time-series baseline, which then allowed us to fit a linear baseline to the spectroscopic light curves of planet b.
-3.2 Relative transmission spectroscopy between planet b and c
-In Figure 1 (lower panel), we present the first relative transmission spectrum between two planets within the same system.
-This relative spectrum is independent of the star’s radius and limb darkening.
-Stellar inhomogeneities are known to contaminate transmission spectra (Désert et al., 2011; Rackham et al., 2019), which has also been highlighted in recent JWST observations(May et al., 2023; Moran et al., 2023; Lim et al., 2023).
-A relative transmission spectrum between planets theoretically bypasses stellar contamination issues, assuming consistent stellar contamination across both transits.
-This enables a better relative comparison of abundances between the planets without precise information on the stellar properties, including metallicity, elemental abundance ratios, and limb-darkening.
-However, in the case of V1298 Tau, an active star, changes in the stellar surface between observation epochs may still introduce contamination to the relative spectrum extracted at different times.
+We ran a free atmospheric chemistry retrieval using the open-source code PetitRadtrans (Mollière et al.
+2019) on the transmission spectrum of planet c.
+We fixed the planet mass and temperature to 17 M⊕ and 950 K, respectively.
+These mass and temperature values for V1298 Tau c are taken from the best fit models to the spectrum from ATMO models, described in Section 3.2.
+We included a grey cloud deck opacity in the free retrieval.
+The posterior distribution from this fit is is shown in Figure 12.
+Our models find water vapor in the atmosphere of planet c, with a lower limit for the water volume mixing ratio ( at 3 confidence level).
+We do not find any constraints on the methane abundance.
+We test the robustness of the water absorption feature indicated by the free retrieval.
+We compared the observed transmission spectrum of V1298 Tau c (Figure 2) to a flat line using a Chi-square test.
+We found that models with atmospheric water absorption are slightly favored (2.5) by comparing the difference in chi-square values between a flat line and the best fit model generated using median values of the fitted parameters from the free retrieval.
+We also tested the impact of fixing the planet mass and isothermal temperature, by including Gaussian priors at 950 K and 17 M⊕ with widths of 10 K and 1 M⊕.
+The retrieved posterior distributions and values for the water and methane abundance agree within 1 to the ones derived from the fixed mass-temperature retrievals, shown in Figure 12.
+Thus, we conclude that with different sets of assumptions, our retrievals suggest the presence of water vapour in the atmosphere of planet c.
+The transit depth uncertainties for planet c are larger by a factor of about 1.5 compared to the spectrum of planet b (Figure 2), adopted from Barat et al.
+(2024).
+In Section 2.2, we discuss the differences in the light curve analysis between planet c and b.
+Since, we cannot ignore the first orbit for the visit of planet c, unlike the case for planet b visit, we cannot apply a common-mode correction to the spectroscopic light curves of planet c.
+Therefore, the higher uncertainty in measured transit depth for V1298 Tau c is likely due to the modelling of each spectroscopic light curve independently using RECTE to model the charge trapping and second order polynomial to model the baseline.
+The common-mode correction could correct for the curvature of the time-series baseline, allowing a linear baseline model as shown in the case of planet b (Barat et al.
+2024).
+3.2 Results from ATMO equilibrium chemistry retrievals
+In this Section, we summarize the results from the equilibrium chemistry atmospheric retrievals using ATMO for both planet b and c.
+The ATMO models are described in Section 2.3.
+Figure 2 shows the best-fit atmospheric models and the observed spectra for both planets.
+The retrieved parameters are enlisted in Table 2, with posterior distributions shown in Figures 13, 14 for planet c and b, respectively.
+We include planet mass as a free parameter for both planets in our retrieval setup.
+We measure a mass (with 1 confidence interval) of 17 M⊕ for planet c.
+The posterior distribution shown in (Figure 13) provides a 3 mass upper limit of 50 M⊕ for planet c.
+The measured mass of planet c from the spectrum is consistent with previous RV mass estimates for this planet: 2010 M⊕ from Sikora et al.
+(2023) and 50 M⊕ at 3 level of confidence (Finociety et al.
+2023).
+Our models find sub-solar/solar atmospheric metallicity for this planet: logZ/Z⊙=-0.16.
+Thus, within 1 confidence interval, the metallicity of planet c is between 0.1-10 solar.
+We do not find any constraint on the cloud deck opacity, the equilibrium temperature or C/O ratio for planet c.
+For planet b, we measure a mass of 8 M⊕ from our free mass retrievals.
+The mass posterior distribution for planet b (Figure 14) shows a 3 upper limit at 20 M⊕.
+The free mass retrieval posterior distribution from Barat et al.
+(2024), which used a different atmospheric model (PetitRadtrans) compared to the present study, found a mass (and 1 confidence interval) of 9 M⊕ and 3 upper limit of 23 M⊕.
+The median value of mass found by Barat et al.
+(2024) is consistent with the present results within 1, and the updated mass upper limit is marginally lower.
+The discrepancy arises because of the non-isothermal T-P profile with high interior temperature and vertical mixing to explain the non-detection of methane for this planet in Barat et al.
+(2024).
+However, our models assume isothermal atmospheres, as a result retrievals favor higher isothermal temperature (800-900 K) compared to Barat et al.
+(2024) (600-700 K) to push the carbon chemistry into methane poor regime (Fortney et al.
+2020).
+Therefore we can conclude that although the mass upper limits slightly differ between Barat et al.
+(2024) and this work, they can be considered consistent given the different model assumptions.
+Our models retrieve a sub-solar atmospheric metallicity (logZ/Z⊙=-2) for planet b.
+This retrieved atmospheric metallicity is consistent within 1 to metallicity reported in Barat et al.
+(2024).
+We found an upper limit to the cloud opacity for this planet: log=-0.6.
+We do not find any constraints on the C/O ratio for planet b.
+We test the impact of keeping the planet mass as a free parameter in our retrievals.
+To assess this, we perform retrievals by fixing the mass of both planets.
+For V1298 Tau c we choose two masses for this test: 10 and 50 M⊕ which are at two ends of the mass posterior distribution (Figure 13) and are also consistent with the mass upper limit from RV studies (Finociety et al.
+2023).
+For planet b, we choose 10 and 23 M⊕.
+Barat et al.
+(2024) has shown that the HST spectrum for V1298 Tau b could be explained with a 10 M⊕ model with a highly sub-solar metallicity and higher cloud opacity, as well as with a relatively clear solar metallicity atmosphere with a planet mass of 23 M⊕.
+We, therefore, choose these two scenarios for this test.
+The posterior distributions for these fixed mass retrievals are shown in Figures 16, 15, with retrieved parameters in Table 2.
+The retrieved atmospheric metallicity from these fixed mass retrievals are consistent within 1 with the free mass retrieval case for both planet b and c (Table 2).
+For planet c, we find that increasing the mass also leads to a small increase in the retrieved atmospheric metallicity, but for all three cases (free mass, 10 M⊕ and 50 M⊕) the retrieved metallicity lies in the range 0.1-10solar for 1 confidence interval.
+For planet b the metallicity is sub-solar for all mass values, however metallicity increases for higher masses.
+Considering the 1 confidence interval, for the three cases for this planet (free mass, 10 M⊕ and 23 M⊕), the atmospheric metallicity lies within 0.01-1 solar metallicity.
+From the atmospheric retrievals on V1298 Tau b and c, we conclude that both these planets have masses in the Neptune or sub-Neptune regime.
+The RV mass for V1298 Tau b from Suárez Mascareño et al.
+(2021) is inconsistent with our measurement, but the planet c mass upper limit from RVs is consistent with our mass measurement for planet c.
+The mass measurements of both planet b and c from this work are consistent with follow-up RV studies (Finociety et al.
+2023; Sikora et al.
+2023).
+The mass we measure for planet b is consistent with previous analysis of its HST transmission spectrum (Barat et al.
+2024).
+Furthermore, TTV analysis of this system (Livingston et al, in prep) also find masses of both these planets in the sub-Neptune mass regime.
+By testing the impact of including planet mass as a free parameter in our retrievals, we conclude that the atmosphere of V1298 Tau b appears to have metallicity that is lower than or equivalent to solar composition.
+For V1298 Tau c, the constraints on the atmospheric metallicity are not as precise compared to planet b; however, the posterior distribution for the metallicity peaks at solar composition with 1 confidence interval between 0.1-10solar.
+Therefore, it is possible that both these planets are young Neptunes or sub-Neptunes in terms of their mass.
+But their atmospheric metallicity is much lower (at least 1 order of magnitude) than what is typically found for known mature Neptune or sub-Neptune planets (e.g.
+see Bean et al.
+2010; Désert et al.
+2011a; Miller-Ricci Kempton et al.
+2012; Morley et al.
+2013; Kreidberg et al.
+2014; Libby-Roberts et al.
+2020a; Bean et al.
+2021).
 | Planet | Parameter | Free mass retrieval (ATMO) | Fixed mass (10 M⊕) | Fixed mass (50 M⊕) |
@@
 | C/O | 0.08 | 0.05 | 0.04 | |
-| Mass [M⊕] | 7.58 | 10 (fixed) | 23 (fixed) |
-3.3 Haze formation in the atmosphere of V1298 Tau planets
-Hazes have been found to be ubiquitous in exoplanet atmospheres (Kempton et al., 2023; Libby-Roberts et al., 2020a; Knutson et al., 2014).
-The temperature of V1298 Tau b and c is expected to be between 600-1000K and the host star is young and active, producing copious amounts of XUV radiation, which are important ingredients for haze formation (Yu et al., 2021).
-We simulated the transmission spectrum of V1298 Tau b and c, including hazes.We assumed tholin like composition for the hazes (Khare et al., 1984).
-The simulations were done following the approach in Kawashima & Ikoma (2019).
-The details of the haze microphysics model are provided in Kawashima & Ikoma (2018).
-We assumed 20M⊕ and 23M⊕ (Barat et al., 2024) for planet c and b respectively.
-For further details see Appendix F.
-The simulated transmission spectrum for V1298 Tau b and c, assuming a 1% haze formation efficiency is shown in Figure 1.
-The effect of hazes in the HST bandpass is to damp the water absorption feature.
-Given the large observable water feature in V1298 Tau b’s spectrum, efficient haze formation can be ruled out for this planet.
-However, atmospheric models with hazes can explain the observed transmission spectrum of planet c.
-If we assume a lower mass for planet c (10M⊕), which is consistent with current estimates from TTV analysis, we would require even higher haze formation rates to explain the observed spectrum of V1298 Tau c.
+| Mass [M⊕] | 8 | 10 (fixed) | 23 (fixed) |
+3.3 Haze formation in the atmosphere of V1298 Tau b and c
+Hazes have been found to be ubiquitous in exoplanet atmospheres (Pont et al.
+2008; Désert et al.
+2011a; Miller-Ricci Kempton et al.
+2012; Morley et al.
+2013; Kreidberg et al.
+2014; Knutson et al.
+2014; Libby-Roberts et al.
+2020a; Kempton et al.
+2023).
+The temperature of V1298 Tau b and c is expected to be between 600-1000K, which has been considered a sweet spot for haze formation (Brande et al.
+2024).
+Furthermore, the host star is young and active, producing copious amounts of UV radiation, which is an important ingredient for haze formation (Kawashima & Ikoma 2018, 2019; Gao et al.
+2020; Yu et al.
+2021).
+We simulated the transmission spectrum of V1298 Tau b and c, including hazes.
+We assumed tholin like composition for the hazes (Khare et al.
+1984).
+The haze microphysics models were simulated using the models described in Kawashima & Ikoma (2018), following the approach of Kawashima & Ikoma (2019).
+First, using a photochemical model the abundance of gaseous species was calculated.
+The XUV spectrum of V1298 Tau was adopted from Duvvuri et al.
+(2023).
+Haze production rates (or efficiency of haze formation) are uncertain; therefore, we assumed a certain fraction of precursor molecules to be dissociated to form haze monomers.
+The haze precursor molecules are assumed to be CH4, HCN and C2H2 in our model.
+We derived vertical density and radius distributions of the haze particles following Kawashima & Ikoma (2018).
+We assumed solar elemental abundance ratio and other parameters such as eddy diffusion coefficient, monomer radius, internal density, and refractive index of haze particles similar to Kreidberg et al.
+(2022).
+We assume 685 K and 979 K as the isothermal temperatures for these models for planet b and c, respectively (Table 1).
+Since the haze models are computationally expensive, we create models for two masses of planet b and c each.
+For planet b, we choose two masses: 10 M⊕ (consistent within 1 to the median of the ATMO retrieval mass posterior distribution for planet b) and 23 M⊕ (3 upper limit from Barat et al.
+(2024)).
+For planet c, we choose 10 M⊕, and 20 M⊕ (consistent within 1 to the median of the ATMO retrieval mass posterior distribution for planet c).
+We choose 10 M⊕ for planet c because from ongoing TTV studies, it is likely that the mass of planet c is in the sub-Neptune regime (Livingston et.
+al in prep).
+In Figure 2 we show the simulated transmission spectra assuming a 1 haze production efficiency for mass of 8 and 17 M⊕ for planet b and c respectively for the ease of comparison with the ATMO models.
+We re-scale the planet c and b models computed with masses of 20 and 10 M⊕ respectively to 17 and 8M⊕ models, assuming that a small change in the planet mass will not significantly affect the haze models.
+The atmospheric signal (which is a relative measurement) is inversely proportional to the planet mass (Heng & Kitzmann 2017), through the scale height.
+Assuming this principle, we re-scale the haze grid models to different mass values.
+While significant haze formation can be ruled out for V1298 Tau b, high haze formation rates could explain the observed spectrum of V1298 Tau c.
+The haze formation efficiency is an uncertain parameter in our models.
+Furthermore, the mass of the planets is also uncertain.
+So, to test the effects of this parameter on the hazy atmospheric models, we create a grid of models at different haze formation efficiency and planet mass.
+In Figure 3 we show the spectra of both planets with different haze formation efficiencies: 10-4 and 10-5 for planet b and 10-5 and 10-2 for planet c.
+For V1298 Tau b, irrespective of the mass, haze formation efficiencies as low as 10-4 can be ruled out as these do not reproduce the large water absorption feature.
+At 8 M⊕, a haze formation efficiency of 10-5 can potentially explain the water absorption feature for planet b, but introduces a strong slope in the continuum which is not seen in the observed spectrum.
+For V1298 Tau c, the haze formation rates cannot be constrained as haze formation efficiencies as low as 10-5 can explain the spectrum.
+However, for the 10 M⊕ for planet c we can rule out low haze formation efficiencies (10-5) at 3.5 level of confidence.
+Therefore, we conclude that V1298 Tau b has a haze formation efficiency lower than 10-5, whereas for V1298 Tau c the atmosphere is likely to have higher haze formation than its sibling planet.
+3.4 Relative transmission spectroscopy between planet b and c
+In Figure 2 (lower panel), we present the first relative transmission spectrum between two planets within the same system.
+By relative transmission spectrum we mean the ratio of transit depths between two planets as a function of wavelength.
+In principle, the relative transmission spectrum is independent of the stellar radii and limb darkening effects, assuming that both planets share similar limb darkening, i.e they transit similar parts of the stellar disk.
+V1298 Tau b and c have impact parameters consistent with each other within 1 (David et al.
+2019b) and relatively low mutual inclination (Johnson et al.
+2022).
+Stellar inhomogeneities are known to contaminate transmission spectra (Désert et al.
+2011a; Rackham et al.
+2019), which has also been highlighted in recent JWST observations(May et al.
+2023; Moran et al.
+2023; Lim et al.
+2023).
+A relative transmission spectrum between multiple planets in the same system, theoretically is independent of stellar contamination, assuming similar stellar contamination across both transits.
+This enables a better relative comparison of abundances between the planets without precise information on the stellar properties, including metallicity, elemental abundance ratios, and limb darkening.
+However, in the case of V1298 Tau, an active star, changes in the stellar surface between observation epochs may still introduce contamination to the relative spectrum extracted at different times.
 4 Discussion
 4.1 Relative orbital distances and stellar density
-In multi-planet systems, it is possible to fix one planet’s semi-major axis based on another’s measured semi-major axis, since they are related to their orbital periods ratio.
+In multi-planet systems, it is possible to fix one planet’s semi-major axis based on another planet’s measured semi-major axis, since they are related to their orbital period ratio.
 However, due to known TTVs in this system, we refrain from fixing the semi-major axis in this analysis.
 The semi-major axis ratio from independent fits to the white light curves of V1298 Tau b and c is 2.080.1 and agrees with the expected value from their known orbital periods.
-Assuming circular orbits, the semi-major axis can be used to estimate stellar density (Sandford & Kipping, 2017).
-From the fits to our white light curves, we estimate the stellar density from the planet b and planet c to be 0.700.01gm/cm3 and 0.670.1gm/cm3.
+Assuming circular orbits, the semi-major axis can be used to estimate stellar density (Sandford & Kipping 2017).
+From the fits to our white light curves, we estimate the stellar density from the planet b and planet c transit light curve to be 0.700.01 gm/cm3 and 0.670.1 gm/cm3.
 These values align with the published stellar density for V1298 Tau (0.680.1 gm/cm3, David et al.
-(2019b)).
-indicating that it is inflation compared to similar-mass main sequence stars.
+(2019a)).
+The relatively low density of V1298 Tau compared to similar mass main-sequence stars indicates that it is inflated and currently undergoing an evolutionary phase, and is likely to end up as a sun-like star as it approaches the main-sequence (David et al.
+2019b).
 4.2 Limb darkening of V1298 Tau
-Modeling stellar limb darkening is crucial for deriving transmission spectra.
-Prior studies often employed fixed limb darkening for main-sequence stars based on theoretical stellar structure models (e.g, see Deming et al., 2013; Sing et al., 2016).
+Modelling stellar limb darkening is crucial for deriving transmission spectra.
+Prior studies often used fixed limb darkening for main-sequence stars based on theoretical stellar structure models (e.g, see Sing et al.
+2016).
 Barat et al.
-(2024) fitted a linear limb darkening coefficient for V1298 Tau, revealing significant differences compared to theoretical models for a main-sequence star with a similar spectral type (K1).
-Higher limb darkening compared to theoretical expectations for mature stars has previously been noted for -Pictoris (Landman et al., 2024).
+(2024) fitted a linear limb darkening coefficient for V1298 Tau, revealing significantly higher limb darkening compared to theoretical models for a main-sequence star with a similar spectral type (K1).
+The limb darkening coefficients have been derived from fitting the transit light curves of V1298 Tau b observed using HST/WFC3 G141.
+These observations had both ingress and egress coverage and allowed precise constraints on the limb darkening model.
+However, for simplicity we choose a linear limb darkening model, rather than other limb darkening models with more free parameters, since HST does not provide continuous time series.
+Higher limb darkening of pre-main sequence stars compared to theoretical models for mature stars has previously been noted for -Pictoris (Landman et al.
+2024).
 The fitted limb darkening coefficients and a theoretical linear limb darkening model is shown in Figure 10.
-This constitutes the first direct measurement of limb darkening on a weak-lined T-Tauri star.
+This is the first direct measurement of limb darkening on a weak-lined T-Tauri star.
 In Section 4.1, we highlighted that V1298 Tau’s density is lower than that of similar-mass main-sequence stars.
-Furthermore, it has a relatively high magnetic field strength of the order of 100-300 Gauss (Finociety et al., 2023).
-High magnetic fields have been proposed as a potential cause for discrepancies between observed and theoretical limb darkening models (Kostogryz et al., 2024).
-Therefore, a combination of structural differences and higher magnetic fields compared to main-sequence stars, could explain the differences we find with the model.
+Furthermore, it has a relatively high magnetic field strength of the order of 100-300 Gauss (Finociety et al.
+2023).
+High magnetic fields have been proposed as a potential cause for discrepancies between observed and theoretical limb darkening models (Kostogryz et al.
+2024).
+Therefore, a combination of structural differences and higher magnetic fields compared to main-sequence stars, could explain the differences we find with the main-sequence limb darkening model.
 While leveraging light curves from multiple planets in the same system could potentially reduce limb darkening coefficient uncertainties, the absence of ingress or egress coverage in the light curves of planet c led us to fix the limb darkening coefficients to the values from Barat et al.
@@
 A comparison of the transmission spectrum of planet c derived using fixed and free limb darkening is shown in Figure 9.
+For the both cases, we assume a linear limb darkening law.
+For the free limb darkening case, we put uniform priors between 0 and 1 on the linear limb darkening coefficient.
 The transit depth uncertainties are higher by 5-10% for each spectroscopic bin in the latter case.
 Residuals in both cases are approximately 1.2 times the expected photon noise, indicating that stellar variability and instrumental systematics from the first orbit are likely to be the dominant sources of noise in our case.
+However, the median subtracted transit depths are consistent within 1 for all spectroscopic bins (Fig 9).
+For the atmospheric retrievals we use the transmission spectrum derived using the fixed limb darkening case.
+In Figure 9 we also show a transmission spectrum for planet c derived by fixing the linear limb darkening coefficient to the EXOCTK model shown in Figure 10.
+We find that the median subtracted transmission spectra for this scenario is consistent within 1 for all spectroscopic bins with the fixed limb darkening case used for the analysis in this paper.
 4.3 Comparative exoplanetology between V1298 Tau b and c
-A comparison between the atmospheres of V1298 Tau b and c, measured relative to atmospheric scale height (Figure 1, upper panel), shows that these two planets could host similar atmospheres.
-Our retrievals show that these planets are likely to have similar masses in the sub-Neptune/super-Earth regime (Table 2).
-Furthermore, we also find that the atmospheric metallicities of both the planets are solar/sub-solar, which is relatively low compared to their mature counterparts which have been found to have metal-rich/hazy (¿solar) atmospheres (Désert et al., 2011; Kempton et al., 2023; Madhusudhan et al., 2023; Libby-Roberts et al., 2020b).
-In this scenario, both the V1298 Tau planets are similar in terms of mass and composition, but their observable atmospheres are distinct from their mature counterparts.
-Atmospheric hazes can be ruled out for V1298 Tau b, given the large atmospheric absorption feature, however for planet c atmospheric hazes can explain the observed spectrum..
-A comparison between the transmission spectra of planet b and c shows that planet c could have a higher haze formation.
-Haze formation has been linked to the UV flux of the host star; a higher UV flux results in higher monomer production rates for organic hazes (Kawashima & Ikoma, 2019).
+In Figure 2 we show a comparison between the atmospheres of V1298 Tau b and c.
+Comparing the spectra in terms of the scale height of both planets (Figure 2, upper panel), shows that these planets could host atmospheres of similar nature.
+To compute the scale height we used the best fit mass from the ATMO posteriors for both planets (Table 2).
+V1298 Tau c has a smaller scale height ( by a factor of 4) compared to V1298 Tau b.
+Atmospheric retrievals show that the masses of both these planets are likely to be in the Neptune/sub-Neptune regime (Section 3.2).
+These models also show that the atmospheric metallicities of both these planets are sub-solar to solar.
+Mature Neptune/sub-Neptune mass planets, are known to have metal-rich/hazy (¿solar) atmospheres (Bean et al.
+2010; Désert et al.
+2011a; Kempton et al.
+2023; Madhusudhan et al.
+2023; Libby-Roberts et al.
+2020b; Bean et al.
+2021).
+Thus, V1298 Tau b and c, which are the two young Neptune/sub-Neptune progenitors, appear to have distinct atmospheres compared to their mature counterparts.
+Efficient haze formation (10) can be ruled out for V1298 Tau b (Figure 3), given the large atmospheric absorption feature.
+However for planet c atmospheric hazes, can explain the observed spectrum.
+In the scenario that the mass of planet c is around 17 M⊕, we cannot constrain the haze formation efficiency for planet c, but for lower masses (10 M⊕), very low haze formation efficiency (10-5) can be ruled out.
+Therefore, planet c could have a higher haze formation efficiency compared to its sibling planet.
+Haze formation has been linked to the UV flux of the host star; a higher UV flux results in higher monomer production rates for organic hazes (Kawashima & Ikoma 2019).
 Higher atmospheric metallicity could also lead to higher haze formation.
@@
 4.4 Comparing the early evolution of V1298 Tau b and c
-We ran evolutionary models for the V1298 Tau planets.
-These models start from the time of disk dispersal, and model the structural and thermal evolution of the planet, including atmospheric mass loss due to photoevaporation (Owen & Wu, 2013).
+We ran evolutionary models for V1298 Tau b and c.
+These models start from the time of disk dispersal, and model the structural and thermal evolution, including atmospheric mass loss due to photoevaporation (Owen & Wu 2013).
+We terminate atmospheric mass loss at 100 Myr.
 We run models assuming both a core-envelope interior structure, as well as with a gradually mixed interior.
@@
 (2022).
-A brief description of the models is provided in Appendix G.
+A brief description of the models is provided in Appendix D.
 Figure 4 shows a possible mass and radius evolution tracks for these planets.
-V1298 Tau c being closer to its host star compared to planet b, is more susceptible to atmospheric mass loss driven by photoevaporation.
-While the final mass and radius depend on photoevaporation efficiency and the star’s activity level (Poppenhaeger et al., 2021), our models show that V1298 Tau c is likely to lose a larger fraction of its current envelope compared to planet b.
-Therefore, even if both planets have similar masses now, their early evolution could lead them to different mass-radius regimes.
+We set up the initial conditions of our models such that these models can match the mass and radius of the planets at the current age (23 Myr).
+For planet c we assume 5.24R⊕ and 17 M⊕ (median value of mass from ATMO retrieval).
+For planet b we assume 9.95 R⊕ and show two models: a 10 M⊕ model which is consistent within 1 to the median value from our ATMO retrievals for planet b, and a 23 M⊕ model which is the 3 mass upper limit from Barat et al.
+(2024).
+We assume that the 10 M⊕ model should be similar to a 8 M⊕ model.
+For this calculation we do not account for the uncertainty in the mass of the planets, therefore, the models should be considered as potential evolutionary tracks.
+Based on these models, V1298 Tau c could have started out with 20 M⊕ with a 25% by mass H/He envelope.
+Currently, it hosts 10% by mass H/He in its envelope and after 100 Myr of evolution it ends up as a Neptune mass planet with 1% by mass H/He envelope.
+After 100 Myr this planet enters into a phase of gradual thermal contraction and ends up with a radius between 2.5-3R⊕ (typical for Neptunes/sub-Neptunes).
+However, for V1298 Tau b, atmospheric mass loss is not as dominant, because it is further away from the host star, thereby, receiving 25% of the stellar flux compared to planet c.
+It starts off with 8.5 M⊕ and 23.5 M⊕ for the two models presented here, and loses less than 1 M⊕ in the course of its evolution.
+However, it is important to note here that Barat et al.
+(2024) report a high internal temperature (400K) for V1298 Tau b from disequilibrium chemistry models of its atmosphere, which has not been considered in our evolution models.
+Higher than expected internal temperature for planet b could lead to a puffed up atmosphere and enhanced atmospheric mass loss compared to our models.
+The mass loss rates are strongly dependent on the current masses of the planets.
+In the scenario that planet b and/or c has mass lower than the estimates used for these calculations the H/He mass fraction in their envelope would be higher, leading to higher mass loss rates.
+In extreme scenarios, these planets could be completely stripped of their primordial atmospheres.
+While the final mass and radius depend on photoevaporation efficiency and the star’s activity level, from these models, we can conclude that due to its proximity to the star, V1298 Tau c is likely to lose a larger fraction of its H/He envelope compared to planet b.
+Therefore, even if both planets have similar masses now, their early evolution could lead them to different mass-radius regimes as well as on different ends of the radius valley.
 Our models assume solar metallicity for the envelope opacity.
 However, the transmission spectrum of V1298 Tau c could also be explained using atmospheric hazes, unlike V1298 Tau b (Figure 2.
-Young planets are expected to cool and contract over time (Kubyshkina et al., 2020; Linder et al., 2019).
-The presence of high altitude hazes can increasing the atmospheric opacity, therefore increasing the cooling and contraction timescale (Lee & Chiang, 2015), leading to higher photoevaporative mass loss.
-Thus, a hazy atmosphere of V1298 Tau c as opposed to a clear atmosphere of V1298 Tau b, could lead to slower cooling and higher mass loss.
-4.5 Comparing the composition of V1298 Tau planet with the exoplanet population
+Young planets are expected to cool and contract over time (Kubyshkina et al.
+2020; Linder et al.
+2019).
+The presence of high altitude hazes can increase the atmospheric opacity, therefore increasing the cooling and contraction timescale (Lee & Chiang 2015), leading to higher photoevaporative mass loss.
+Thus, a hazy atmosphere of V1298 Tau c as opposed to a clear atmosphere of V1298 Tau b, could also lead to slower cooling and higher mass loss rates for planet c.
+4.5 Comparing the atmospheres of V1298 Tau b and c with the exoplanet population
 Figure 5 shows V1298 Tau b and c on the mass-metallicity and metallicity-orbital distance planes.
-The observed low atmospheric metallicity for these planets deviates from the mass-metallicity trend predicted by core-accretion models (Pollack et al., 1996; Bodenheimer & Pollack, 1986) and reported for exoplanets (Welbanks et al., 2019; Thorngren & Fortney, 2019).
-This could potentially be explained by ongoing evolutionary processes that transform their current metal-poor atmospheres into the metal-rich atmospheres seen in mature sub-Neptune/super-Earth planets (Désert et al., 2011; Kempton et al., 2023; Libby-Roberts et al., 2020a; Madhusudhan et al., 2023).
-We compare the V1298 Tau system with the 30 Myr old HR 8799 system (Marois et al., 2010) with four directly imaged planets (HR8799 b, c , d and e) with masses 7, 10, 10, 10 (Marois et al., 2010; GRAVITY Collaboration et al., 2019).
+The observed low atmospheric metallicity (solar/sub-solar) for these planets, deviates from the mass-metallicity trend predicted by core-accretion models (Bodenheimer & Pollack 1986; Pollack et al.
+1996) and reported for exoplanets (Welbanks et al.
+2019; Thorngren & Fortney 2019).
+The atmospheric metallicities of planet b and c appears to be at least one order of magnitude lower compared to mature Neptune/sub-Neptunes.
+This could potentially be explained by ongoing evolutionary processes that transform their current metal-poor atmospheres of V1298 Tau b and c into the metal-rich atmospheres, as found in known mature Neptune/sub-Neptune planets (Désert et al.
+2011a; Miller-Ricci Kempton et al.
+2012; Morley et al.
+2013; Libby-Roberts et al.
+2020a; Bean et al.
+2021; Kempton et al.
+2023; Madhusudhan et al.
+2023).
+One such possible mechanism has been suggested, where Neptune/sub-Neptunes are naturally born with metallicity gradients in their envelope (Ormel et al.
+2021).
+Atmospheric mass loss during the early evolutionary phase removes the upper layers of the atmosphere, revealing the inner metal-rich layers (Fortney et al.
+2013; Vazan et al.
+2022).
+Thus, this mechanism could reconcile the the two Neptune/sub-Neptune progenitors, V1298 Tau b and c with their mature counterparts.
+We compare the V1298 Tau system with the 30 Myr old HR 8799 system (Marois et al.
+2010) with four directly imaged planets (HR8799 b, c , d and e) with masses 7, 10, 10, 10 MJ, respectively (Marois et al.
+2010; GRAVITY Collaboration et al.
+2019).
 Nasedkin et al.
 (2024) using ground based spectra from the VLTI/GRAVITY program, report super-stellar(100 times stellar) atmospheric metallicity for these planets.
-Assuming a -0.65 dex metallicity for HR8799 (Swastik et al., 2021) we convert the retrieved atmospheric metallicity to stellar unit.
-The HR 8799 planets exhibit higher atmospheric metallicity compared to the mass-metallicity trend (Thorngren & Fortney, 2019) predicted by core-accretion formation, while V1298 Tau planets fall below this trend.
-HR 8799 is a system of widely separated massive gas giants, whereas V1298 Tau likely consists of compact, low-mass sub-Neptune/super-Earth progenitors.
+Assuming a -0.65 dex metallicity for HR8799 (Swastik et al.
+2021) we convert the retrieved atmospheric metallicity to stellar unit.
+The HR 8799 planets exhibit higher atmospheric metallicity compared to the mass-metallicity trend (Thorngren & Fortney 2019) predicted by core-accretion formation, while V1298 Tau planets fall below this trend.
+HR 8799 is a system of widely separated massive gas giants, whereas V1298 Tau likely consists of compact, low-mass Neptune/sub-Neptune progenitors.
 Both systems likely host primordial atmospheres.
-The different mass regimes suggest HR 8799 planets underwent runaway gas accretion or gravitational collapse, unlike V1298 Tau planets.
-Theories propose close-in sub-Neptunes and super-Earths formed in situ within the water ice-line in dusty or depleting disks to prevent runaway gas accretion (Lee et al., 2014; Lee & Chiang, 2016).
-Wide orbit gas giants likely formed beyond the ice lines, allowing for runaway accretion (Pollack et al., 1996) or gravitational instability (Boss, 1997; Dodson-Robinson et al., 2009), but these mechanisms alone do not explain HR 8799’s high metallicity, necessitating the accretion of volatile-rich solids (Nasedkin et al., 2024).
-Formation location relative to ice lines could lead to differences in volatile enrichment, as beyond the ice-line, the disk contains volatile-rich solids, enhancing the metal content of primordial atmospheres (Öberg et al., 2011; Bitsch et al., 2019).
+The different mass regimes suggest HR 8799 planets underwent runaway gas accretion (Pollack et al.
+1996) or gravitational collapse (Boss 1997), unlike V1298 Tau planets.
+Formation theories of Neptune/sub-Neptunes suggest that these planets are born in situ within the water ice-line in dusty or depleting disks to prevent runaway gas accretion (Lee et al.
+2014; Lee & Chiang 2016).
+Wide orbit gas giants likely formed beyond the ice lines, allowing for runaway accretion (Pollack et al.
+1996) or gravitational instability (Boss 1997; Dodson-Robinson et al.
+2009), but these mechanisms alone do not explain HR 8799’s high metallicity, necessitating the accretion of volatile-rich solids (Nasedkin et al.
+2024).
+Formation location relative to ice lines could lead to differences in volatile enrichment, as beyond the ice-line, the disk contains volatile-rich solids, enhancing the metal content of primordial atmospheres (Öberg et al.
+2011; Bitsch et al.
+2019).
 4.6 Formation and evolution of multi-planet systems
-The mass of the V1298 Tau planets indicates they are likely sub-Neptune/super-Earth progenitors, which are in a near mean-motion resonance (1:2:3).
-Atmospheric retrievals suggest these planets could have low metallicity atmospheres.
-Population synthesis studies (Mordasini, 2018; Benz et al., 2014) show sub-Neptunes and super-Earths often form in multi-planet systems in mean motion resonance (Lambrechts et al., 2019).
+The mass of V1298 Tau b and c indicate that they are likely Neptune/sub-Neptune progenitors, which are in a near mean-motion resonance (1:2:3).
+Atmospheric retrievals suggest both these planets could have relatively low metallicity atmospheres.
+Population synthesis studies (Benz et al.
+2014; Mordasini 2018) show that low mass planets, often form in multi-planet systems in mean motion resonance (Lambrechts et al.
+2019).
 Emsenhuber et al.
 (2023) classified systems with multiple low-mass planets into two types: Class I: inner dry/rocky planets formed in situ with outer icy sub-Neptunes migrating inwards, and Class II: multiple sub-Neptunes forming beyond the water ice line and migrating inwards.
-In the first scenario, inner planets have volatile-poor atmospheres, whereas, in the second, all planets are volatile-rich.
+In the first scenario, inner planets have metal-poor atmospheres, whereas, in the second, all planets are metal-rich.
 The relatively low atmospheric metallicity of V1298 Tau b and c suggests a formation similar to class I planets .
-However, these young planets’ atmospheric composition may evolve.
-In the pebble accretion scenario, a metallicity gradient forms in the protoplanet’s envelope (Ormel et al., 2021; Bodenheimer et al., 2018), with metal-rich layers hidden deep inside.
-Over time, atmospheric mass loss can strip volatile-poor upper layers, enriching atmospheric metallicity (Vazan et al., 2022), aligning them with mature sub-Neptune/super-Earth population.
+However, these young planets’ atmospheric metallicity may evolve.
+In the pebble accretion scenario, a metallicity gradient forms in the protoplanet’s envelope (Bodenheimer et al.
+2018; Ormel et al.
+2021), with metal-rich layers hidden deep inside.
+Over time, atmospheric mass loss can strip volatile-poor upper layers, revealing deeper metal-rich atmospheric layers(Fortney et al.
+2013; Vazan et al.
+2022), aligning them with mature sub-Neptune/super-Earth population.
 In this case, the V1298 Tau system could also have a formation similar to the class II systems.
@@
 Class I systems have inner volatile-poor planets, while class II systems have planets with similar compositions.
-Due to its proximity to the star, V1298 Tau c is likely to lose more of its primordial envelope compared to V1298 Tau b 2.
-Thus, V1298 Tau b and c might have formed like class II systems, with volatiles hidden deep inside, but atmospheric mass loss caused planet c’s composition to evolve faster than planet b, explaining their difference in their atmospheric composition.
+Due to its proximity to the star, V1298 Tau c is likely to lose more of its primordial envelope compared to V1298 Tau b 4.
+Thus, V1298 Tau b and c might have formed like class II systems, with volatiles hidden deep inside, but atmospheric mass loss caused planet c’s metallicity to evolve faster than planet b.
 5 Conclusion
-In this paper, we perform the first comparative study of atmospheres of multiple planets in the same transiting system.
-Comparing the transmission spectra of the young V1298 Tau b and c planets reveals that planet b has a large, haze-free envelope, while the presence of hazes in planet c cannot be ruled out.
-Atmospheric models suggest similar masses for both planets in the sub-Neptune/super-Earth regime.
-These models also find relatively low atmospheric metallicities for both these young planets compared to their matured counterparts, indicating they are sub-Neptune/super-Earth progenitors.
-Alternatively, if planet c has a hazy atmosphere, higher haze formation in planet c compared to planet b could be due to differences in composition, temperature, and incoming stellar UV flux.
-The higher envelope opacity of planet c would affect its cooling rate and evolution differently than for planet b.
-6 Acknowledgements
+In this paper, we perform the first comparative study of atmospheres of multiple planets in the same transiting multi-planet system.
+We compare the transmission spectra of two exoplanets in the young (20-30 Myr old) V1298 Tau system: V1298 Tau b and c.
+The transmission spectrum of V1298 Tau c shows an absorption feature that we interpret as due to water vapour (2.5), whereas, the spectra of its sibling, V1298 Tau b (adopted from Barat et al.
+(2024)), shows a clear atmosphere with a large scale height and a water absorption feature at 5 level of significance (Figure 2).
+We measure the masses of both planets from the transmission spectra ( M⊕ for planet b and M⊕ for planet c), together with the chemical abundances using a Bayesian chemical equilibrium retrieval framework using the ATMO atmospheric model.
+The retrieved mass for planet c is consistent with previous RV estimates (Suárez Mascareño et al.
+2021; Sikora et al.
+2023; Finociety et al.
+2023).
+Our mass determination from the atmospheric scale heights also agree with ongoing TTV studies (Livingston et al, in prep).
+Our atmospheric models find sub-solar/solar atmospheric metallicity (logZ/Z⊙=-2.04 for planet b and logZ/Z⊙= -0.16 for planet c) for both these planets (Table 2), from free planet masses as well as fixed masses in the atmospheric retrievals.
+Compared to known mature Neptune/sub-Neptunes, which have been found to have metal-rich (100solar) and/or hazy atmospheres, the atmospheres of the two young Neptune/sub-Neptune progenitors (planet b and c), can be considered metal-poor by at least one order of magnitude.
+This hints at possible ongoing early evolutionary mechanisms which are expected to enhance the atmospheric metallicity of these planets, such that they are reconciled with their mature counterparts after evolution.
+Alternatively, the observed spectrum of V1298 Tau c could also be explained by atmospheric hazes.
+The large observed scale height of V1298 Tau b rules out high haze formation efficiency (, Figure 3) for this planet.
+Higher haze formation in planet c compared to planet b could be due to differences in bulk composition, temperature-pressure profile, and incoming stellar UV flux, which are known to be important factors affecting haze formation in H/He rich atmospheres (Kawashima & Ikoma 2018; Gao & Zhang 2020).
+In this scenario, the higher envelope opacity of planet c due to hazes could delay its cooling rate and compared to planet b, which has a clearer atmosphere.
+Evolutionary models suggest that planet c could lose a large fraction of its current H/He envelope (10% by mass) and end up with % by the end of it’s evolutionary phase.
+In contrast, our models applied to planet b, which receives 25% of the stellar flux received by planet c, do not predict significant mass loss and could end up retaining a substantial H/He envelope after 100 Myr of evolution.
+Acknowledgements.
 J.M.D acknowledges support from the Amsterdam Academic Alliance (AAA) Program, and the European Research Council (ERC) European Union’s Horizon 2020 research and innovation program (grant agreement no.
@@
 Y.K.
-acknowledges the support from JSPS KAKENHI Grant Numbers JP21K13984 and 22H05150.
-7 Softwares used
-numpy, scipy, matplotlib, emcee, lmfit, pandas, ATMO, PetitRadtrans
+acknowledges the support from JSPS KAKENHI Grant Numbers 21K13984, 22H05150, and 23H01224.
+We acknowledge publicly available open source softwares which have been used for this work: numpy, scipy, matplotlib, emcee, lmfit, pandas, ATMO, PetitRadtransReferences
+- Ahrer et al. (2023) Ahrer, E.-M., Stevenson, K. B., Mansfield, M., et al. 2023, Nature, 614, 653
+- Alderson et al. (2023) Alderson, L., Wakeford, H. R., Alam, M. K., et al. 2023, Nature, 614, 664
+- Arcangeli et al. (2018) Arcangeli, J., Désert, J.-M., Line, M. R., et al. 2018, The Astrophysical Journal, 855, L30
+- Arcangeli et al. (2019) Arcangeli, J., Désert, J.-M., Parmentier, V., et al. 2019, A&A, 625, A136
+- Arcangeli et al. (2021) Arcangeli, J., Désert, J. M., Parmentier, V., Tsai, S. M., & Stevenson, K. B. 2021, A&A, 646, A94
+- August et al. (2023) August, P. C., Bean, J. L., Zhang, M., et al. 2023, The Astrophysical Journal Letters, 953, L24
+- Baeyens et al. (2022) Baeyens, R., Konings, T., Venot, O., Carone, L., & Decin, L. 2022, MNRAS, 512, 4877
+- Barat et al. (2024) Barat, S., Désert, J.-M., Vazan, A., et al. 2024, Nature Astronomy [arXiv:2312.16924]
+- Barclay et al. (2021) Barclay, T., Kostov, V. B., Colón, K. D., et al. 2021, AJ, 162, 300
+- Baxter et al. (2020) Baxter, C., Désert, J.-M., Parmentier, V., et al. 2020, A&A, 639, A36
+- Baxter et al. (2021) Baxter, C., Désert, J.-M., Tsai, S.-M., et al. 2021, A&A, 648, A127
+- Bean et al. (2010) Bean, J. L., Miller-Ricci Kempton, E., & Homeier, D. 2010, Nature, 468, 669
+- Bean et al. (2021) Bean, J. L., Raymond, S. N., & Owen, J. E. 2021, Journal of Geophysical Research (Planets), 126, e06639
+- Bell et al. (2023) Bell, T. J., Welbanks, L., Schlawin, E., et al. 2023, Nature, 623, 709
+- Benneke et al. (2024) Benneke, B., Roy, P.-A., Coulombe, L.-P., et al. 2024, arXiv e-prints, arXiv:2403.03325
+- Benz et al. (2014) Benz, W., Ida, S., Alibert, Y., Lin, D., & Mordasini, C. 2014, in Protostars and Planets VI, ed. H. Beuther, R. S. Klessen, C. P. Dullemond, & T. Henning, 691–713
+- Berta et al. (2012) Berta, Z. K., Charbonneau, D., Désert, J.-M., et al. 2012, ApJ, 747, 35
+- Bitsch et al. (2019) Bitsch, B., Raymond, S. N., & Izidoro, A. 2019, A&A, 624, A109
+- Blunt et al. (2023) Blunt, S., Carvalho, A., David, T. J., et al. 2023, AJ, 166, 62
+- Bodenheimer & Pollack (1986) Bodenheimer, P. & Pollack, J. B. 1986, Icarus, 67, 391
+- Bodenheimer et al. (2018) Bodenheimer, P., Stevenson, D. J., Lissauer, J. J., & D’Angelo, G. 2018, ApJ, 868, 138
+- Boss (1997) Boss, A. P. 1997, Science, 276, 1836
+- Bourque et al. (2021) Bourque, M., Espinoza, N., Filippazzo, J., et al. 2021, The Exoplanet Characterization Toolkit (ExoCTK)
+- Brande et al. (2024) Brande, J., Crossfield, I. J. M., Kreidberg, L., et al. 2024, The Astrophysical Journal Letters, 961, L23
+- Brems et al. (2019) Brems, S. S., Kürster, M., Trifonov, T., Reffert, S., & Quirrenbach, A. 2019, A&A, 632, A37
+- Carter et al. (2020) Carter, A. L., Nikolov, N., Sing, D. K., et al. 2020, MNRAS, 494, 5449
+- Coulombe et al. (2023) Coulombe, L.-P., Benneke, B., Challener, R., et al. 2023, Nature, 620, 292
+- Damasso et al. (2023) Damasso, M., Scandariato, G., Nascimbeni, V., et al. 2023, A&A, 680, A8
+- David et al. (2019a) David, T. J., Cody, A. M., Hedges, C. L., et al. 2019a, AJ, 158, 79
+- David et al. (2016) David, T. J., Hillenbrand, L. A., Petigura, E. A., et al. 2016, Nature, 534, 658
+- David et al. (2019b) David, T. J., Petigura, E. A., Luger, R., et al. 2019b, ApJ, 885, L12
+- Deming et al. (2013) Deming, D., Wilkins, A., McCullough, P., et al. 2013, ApJ, 774, 95
+- Désert et al. (2011a) Désert, J.-M., Bean, J., Miller-Ricci Kempton, E., et al. 2011a, ApJ, 731, L40
+- Désert et al. (2011b) Désert, J.-M., Charbonneau, D., Demory, B.-O., et al. 2011b, ApJS, 197, 14
+- Dodson-Robinson et al. (2009) Dodson-Robinson, S. E., Veras, D., Ford, E. B., & Beichman, C. A. 2009, The Astrophysical Journal, 707, 79
+- Drummond et al. (2016) Drummond, B., Tremblin, P., Baraffe, I., et al. 2016, A&A, 594, A69
+- Duvvuri et al. (2023) Duvvuri, G. M., Cauley, P. W., Aguirre, F. C., et al. 2023, The Astronomical Journal, 166, 196
+- Emsenhuber et al. (2023) Emsenhuber, A., Mordasini, C., & Burn, R. 2023, European Physical Journal Plus, 138, 181
+- Erkaev et al. (2007) Erkaev, N. V., Kulikov, Y. N., Lammer, H., et al. 2007, A&A, 472, 329
+- Evans et al. (2018) Evans, T. M., Sing, D. K., Goyal, J. M., et al. 2018, AJ, 156, 283
+- Feinstein et al. (2022) Feinstein, A. D., David, T. J., Montet, B. T., et al. 2022, The Astrophysical Journal Letters, 925, L2
+- Feinstein et al. (2021) Feinstein, A. D., Montet, B. T., Johnson, M. C., et al. 2021, AJ, 162, 213
+- Feinstein et al. (2023) Feinstein, A. D., Radica, M., Welbanks, L., et al. 2023, Nature, 614, 670
+- Finociety et al. (2023) Finociety, B., Donati, J. F., Cristofari, P. I., et al. 2023, MNRAS, 526, 4627
+- Foreman-Mackey et al. (2013) Foreman-Mackey, D., Hogg, D. W., Lang, D., & Goodman, J. 2013, PASP, 125, 306
+- Fortney et al. (2013) Fortney, J. J., Mordasini, C., Nettelmann, N., et al. 2013, 775, 80
+- Fortney et al. (2020) Fortney, J. J., Visscher, C., Marley, M. S., et al. 2020, AJ, 160, 288
+- Fu et al. (2017) Fu, G., Deming, D., Knutson, H., et al. 2017, The Astrophysical Journal Letters, 847, L22
+- Fulton & Petigura (2018) Fulton, B. J. & Petigura, E. A. 2018, AJ, 156, 264
+- Fulton et al. (2017) Fulton, B. J., Petigura, E. A., Howard, A. W., et al. 2017, The Astronomical Journal, 154, 109
+- Gagnebin et al. (2024) Gagnebin, A., Mukherjee, S., Fortney, J. J., & Batalha, N. E. 2024, ApJ, 969, 86
+- Gao et al. (2020) Gao, P., Thorngren, D. P., Lee, E. K. H., et al. 2020, Nature Astronomy, 4, 951
+- Gao & Zhang (2020) Gao, P. & Zhang, X. 2020, ApJ, 890, 93
+- Garhart et al. (2020) Garhart, E., Deming, D., Mandell, A., et al. 2020, AJ, 159, 137
+- Ginzburg et al. (2018) Ginzburg, S., Schlichting, H. E., & Sari, R. 2018, MNRAS, 476, 759
+- Goyal et al. (2018) Goyal, J. M., Mayne, N., Sing, D. K., et al. 2018, MNRAS, 474, 5158
+- Goyal et al. (2019) Goyal, J. M., Wakeford, H. R., Mayne, N. J., et al. 2019, MNRAS, 482, 4503
+- Grant et al. (2023) Grant, D., Lewis, N. K., Wakeford, H. R., et al. 2023, The Astrophysical Journal Letters, 956, L32
+- GRAVITY Collaboration et al. (2019) GRAVITY Collaboration, Lacour, S., Nowak, M., et al. 2019, A&A, 623, L11
+- Gupta & Schlichting (2019) Gupta, A. & Schlichting, H. E. 2019, Monthly Notices of the Royal Astronomical Society, 487, 24
+- Heng & Kitzmann (2017) Heng, K. & Kitzmann, D. 2017, MNRAS, 470, 2972
+- Horne (1986) Horne, K. 1986, PASP, 98, 609
+- Jacobs et al. (2022) Jacobs, B., Désert, J. M., Pino, L., et al. 2022, A&A, 668, L1
+- Johnson et al. (2022) Johnson, M. C., David, T. J., Petigura, E. A., et al. 2022, AJ, 163, 247
+- Kajtazi et al. (2023) Kajtazi, K., Petit, A. C., & Johansen, A. 2023, A&A, 669, A44
+- Kawashima & Ikoma (2018) Kawashima, Y. & Ikoma, M. 2018, The Astrophysical Journal, 853, 7
+- Kawashima & Ikoma (2019) Kawashima, Y. & Ikoma, M. 2019, ApJ, 877, 109
+- Keating et al. (2019) Keating, D., Cowan, N. B., & Dang, L. 2019, Nature Astronomy, 3, 1092
+- Kempton et al. (2023) Kempton, E. M. R., Zhang, M., Bean, J. L., et al. 2023, Nature, 620, 67
+- Khare et al. (1984) Khare, B. N., Sagan, C., Arakawa, E. T., et al. 1984, Icarus, 60, 127
+- Knutson et al. (2014) Knutson, H. A., Benneke, B., Deming, D., & Homeier, D. 2014, Nature, 505, 66
+- Koen (2016) Koen, C. 2016, MNRAS, 463, 4383
+- Kostogryz et al. (2024) Kostogryz, N. M., Shapiro, A. I., Witzke, V., et al. 2024, arXiv e-prints, arXiv:2403.00118
+- Kreidberg (2015) Kreidberg, L. 2015, PASP, 127, 1161
+- Kreidberg et al. (2014) Kreidberg, L., Bean, J. L., Désert, J.-M., et al. 2014, Nature, 505, 69
+- Kreidberg et al. (2022) Kreidberg, L., Mollière, P., Crossfield, I. J. M., et al. 2022, AJ, 164, 124
+- Kubyshkina et al. (2020) Kubyshkina, D., Vidotto, A. A., Fossati, L., & Farrell, E. 2020, MNRAS, 499, 77
+- Lambrechts et al. (2019) Lambrechts, M., Morbidelli, A., Jacobson, S. A., et al. 2019, A&A, 627, A83
+- Landman et al. (2024) Landman, R., Stolker, T., Snellen, I. A. G., et al. 2024, A&A, 682, A48
+- Lee & Chiang (2015) Lee, E. J. & Chiang, E. 2015, The Astrophysical Journal, 811, 41
+- Lee & Chiang (2016) Lee, E. J. & Chiang, E. 2016, ApJ, 817, 90
+- Lee et al. (2014) Lee, E. J., Chiang, E., & Ormel, C. W. 2014, ApJ, 797, 95
+- Lewis et al. (2020) Lewis, N. K., Wakeford, H. R., MacDonald, R. J., et al. 2020, ApJ, 902, L19
+- Libby-Roberts et al. (2020a) Libby-Roberts, J. E., Berta-Thompson, Z. K., Désert, J.-M., et al. 2020a, AJ, 159, 57
+- Libby-Roberts et al. (2020b) Libby-Roberts, J. E., Berta-Thompson, Z. K., Désert, J.-M., et al. 2020b, AJ, 159, 57
+- Lim et al. (2023) Lim, O., Benneke, B., Doyon, R., et al. 2023, ApJ, 955, L22
+- Linder et al. (2019) Linder, E. F., Mordasini, C., Mollière, P., et al. 2019, A&A, 623, A85
+- Luhman (2018) Luhman, K. L. 2018, AJ, 156, 271
+- Madhusudhan et al. (2023) Madhusudhan, N., Sarkar, S., Constantinou, S., et al. 2023, ApJ, 956, L13
+- Mann et al. (2016) Mann, A. W., Newton, E. R., Rizzuto, A. C., et al. 2016, The Astronomical Journal, 152, 61
+- Mann et al. (2022) Mann, A. W., Wood, M. L., Schmidt, S. P., et al. 2022, AJ, 163, 156
+- Marois et al. (2010) Marois, C., Zuckerman, B., Konopacky, Q. M., Macintosh, B., & Barman, T. 2010, Nature, 468, 1080
+- May et al. (2023) May, E. M., MacDonald, R. J., Bennett, K. A., et al. 2023, ApJ, 959, L9
+- Miller-Ricci Kempton et al. (2012) Miller-Ricci Kempton, E., Zahnle, K., & Fortney, J. J. 2012, ApJ, 745, 3
+- Mollière et al. (2019) Mollière, P., Wardenier, J. P., van Boekel, R., et al. 2019, A&A, 627, A67
+- Moran et al. (2023) Moran, S. E., Stevenson, K. B., Sing, D. K., et al. 2023, ApJ, 948, L11
+- Mordasini (2018) Mordasini, C. 2018, in Handbook of Exoplanets, ed. H. J. Deeg & J. A. Belmonte, 143
+- Morley et al. (2013) Morley, C. V., Fortney, J. J., Kempton, E. M. R., et al. 2013, ApJ, 775, 33
+- Nasedkin et al. (2024) Nasedkin, E., Mollière, P., Lacour, S., et al. 2024, A&A, 687, A298
+- Öberg et al. (2011) Öberg, K. I., Murray-Clay, R., & Bergin, E. A. 2011, ApJ, 743, L16
+- Oh et al. (2017) Oh, S., Price-Whelan, A. M., Hogg, D. W., Morton, T. D., & Spergel, D. N. 2017, VizieR Online Data Catalog, J/AJ/153/257
+- Ormel et al. (2021) Ormel, C. W., Vazan, A., & Brouwers, M. G. 2021, A&A, 647, A175
+- Owen & Wu (2013) Owen, J. E. & Wu, Y. 2013, ApJ, 775, 105
+- Owen & Wu (2017) Owen, J. E. & Wu, Y. 2017, The Astrophysical Journal, 847, 29
+- Pollack et al. (1996) Pollack, J. B., Hubickyj, O., Bodenheimer, P., et al. 1996, Icarus, 124, 62
+- Pont et al. (2008) Pont, F., Knutson, H., Gilliland, R. L., Moutou, C., & Charbonneau, D. 2008, MNRAS, 385, 109
+- Rackham et al. (2019) Rackham, B. V., Apai, D., & Giampapa, M. S. 2019, AJ, 157, 96
+- Radica et al. (2023) Radica, M., Welbanks, L., Espinoza, N., et al. 2023, MNRAS, 524, 835
+- Ranjan et al. (2014) Ranjan, S., Charbonneau, D., Désert, J.-M., et al. 2014, ApJ, 785, 148
+- Rathcke et al. (2021) Rathcke, A. D., MacDonald, R. J., Barstow, J. K., et al. 2021, AJ, 162, 138
+- Rogers & Owen (2021) Rogers, J. G. & Owen, J. E. 2021, MNRAS, 503, 1526
+- Rustamkulov et al. (2023) Rustamkulov, Z., Sing, D. K., Mukherjee, S., et al. 2023, Nature, 614, 659
+- Sandford & Kipping (2017) Sandford, E. & Kipping, D. 2017, The Astronomical Journal, 154, 228
+- Shulyak et al. (2020) Shulyak, D., Lara, L. M., Rengel, M., & Nèmec, N. E. 2020, A&A, 639, A48
+- Sikora et al. (2023) Sikora, J., Rowe, J., Barat, S., et al. 2023, AJ, 165, 250
+- Sing et al. (2016) Sing, D. K., Fortney, J. J., Nikolov, N., et al. 2016, Nature, 529, 59
+- Sing et al. (2024) Sing, D. K., Rustamkulov, Z., Thorngren, D. P., et al. 2024, Nature, 630, 831
+- Speagle (2020) Speagle, J. S. 2020, MNRAS, 493, 3132
+- Suárez Mascareño et al. (2021) Suárez Mascareño, A., Damasso, M., Lodieu, N., et al. 2021, Nature Astronomy [arXiv:2111.09193]
+- Swastik et al. (2021) Swastik, C., Banyal, R. K., Narang, M., et al. 2021, AJ, 161, 114
+- Szabó & Kiss (2011) Szabó, G. M. & Kiss, L. L. 2011, ApJ, 727, L44
+- Terquem & Papaloizou (2018) Terquem, C. & Papaloizou, J. C. B. 2018, Monthly Notices of the Royal Astronomical Society, 482, 530
+- Thorngren & Fortney (2019) Thorngren, D. & Fortney, J. J. 2019, ApJ, 874, L31
+- Tremblin et al. (2015) Tremblin, P., Amundsen, D. S., Mourier, P., et al. 2015, ApJ, 804, L17
+- Tsai et al. (2023) Tsai, S.-M., Lee, E. K. H., Powell, D., et al. 2023, Nature, 617, 483
+- Tsiaras et al. (2016) Tsiaras, A., Waldmann, I. P., Rocchetto, M., et al. 2016, The Astrophysical Journal, 832, 202
+- Van Eylen et al. (2018) Van Eylen, V., Agentoft, C., Lundkvist, M. S., et al. 2018, MNRAS, 479, 4786
+- Vazan et al. (2022) Vazan, A., Sari, R., & Kessel, R. 2022, The Astrophysical Journal, 926, 150
+- Venot & Agúndez (2015) Venot, O. & Agúndez, M. 2015, Experimental Astronomy, 40, 469
+- Wakeford & Dalba (2020) Wakeford, H. R. & Dalba, P. A. 2020, Philosophical Transactions of the Royal Society of London Series A, 378, 20200054
+- Wakeford et al. (2013) Wakeford, H. R., Sing, D. K., Deming, D., et al. 2013, Monthly Notices of the Royal Astronomical Society, 435, 3481
+- Wakeford et al. (2016) Wakeford, H. R., Sing, D. K., Evans, T., Deming, D., & Mandell, A. 2016, ApJ, 819, 10
+- Welbanks et al. (2019) Welbanks, L., Madhusudhan, N., Allard, N. F., et al. 2019, ApJ, 887, L20
+- Xue et al. (2024) Xue, Q., Bean, J. L., Zhang, M., et al. 2024, ApJ, 963, L5
+- Yu et al. (2021) Yu, X., He, C., Zhang, X., et al. 2021, Nature Astronomy, 5, 822
+- Zhou et al. (2020) Zhou, G., Quinn, S. N., Irwin, J., et al. 2020, The Astronomical Journal, 161, 2
+- Zhou et al. (2017) Zhou, Y., Apai, D., Lew, B. W. P., & Schneider, G. 2017, The Astronomical Journal, 153, 243
 Appendix A Description of raw data reduction
+We use the HST/WFC3 data reduction pipeline which has been presented in previous studies (Arcangeli et al.
+2018, 2019, 2021; Jacobs et al.
+2022; Barat et al.
+2024).
+We do not change anything in the pipeline for the analysis of this paper.
 We start our analysis from the ‘ima.fits’ files downloaded from MAST archive.
@@
 We create sub-exposures by subtracting consecutive non-destructive reads for each exposure which are reduced separately.
-This technique has been widely applied in past observations with HST (e.g, see Kreidberg et al.
-2014; Arcangeli et al.
-2018; Jacobs et al.
-2022).
+This technique has been widely applied in past observations with HST.
 We apply a wavelength dependent flat-field correction and flag bad pixels with data quality DQ=4, 32, or 512 by calwf3.
@@
 Appendix B Light curve fitting
-HST WFC3 light curves often exhibit strong systematics like charge-trapping ramps and visit-long slopes (Kreidberg et al.
-2014; Deming et al.
-2013; Tsiaras et al.
+HST WFC3 light curves often exhibit strong systematics, such as ‘hook-like’ ramps and visit-long slopes (Berta et al.
+2012; Deming et al.
+2013; Ranjan et al.
+2014; Tsiaras et al.
 2016; Wakeford et al.
 2016).
-Common practice involves discarding the first orbit due to significantly larger systematics (See Figure 4) compared to the rest of the visit (e.g see, Wakeford et al.
+Common practice involves discarding the first orbit due to significantly larger systematics (See Figure 1) compared to the rest of the visit (e.g.
+see, Wakeford et al.
 2013; Arcangeli et al.
@@
 2022; Barat et al.
+2024) from the light curve fitting.
+For the visit of planet b, the first orbit had been discarded (Barat et al.
 2024).
-However, in this case, we could not remove from the analysis because the transit occurred two hours earlier than expected from the linear ephemeris (Feinstein et al.
+However, for the visit of planet c, we could not remove the first orbit of the visit from the analysis because the transit occurred two hours earlier than expected from the linear ephemeris (Feinstein et al.
 2022) due to large TTVs and we could only secure one orbit prior to ingress.
 The first-orbit systematics are strongly wavelength dependent (Zhou et al.
-2020, 2017) and cannot be corrected for using common-mode approaches, which had been previously applied to similar observations of V1298 tau b (Barat et al.
+2020, 2017).
+Therefore, the spectroscopic light curves cannot be de-trended using a common-mode approach, which had been previously applied to similar observations of V1298 tau b (Barat et al.
 2024).
 Therefore, to handle these systematics, we used the physically motivated RECTE charge-trapping model (Zhou et al.
-2017), which fits the first orbit ramp better than other methods (Berta et al.
-2012; Kreidberg et al.
-2014).
+2017), which fits the first orbit ramp better than other known methods (e.g, Berta et al.
+2012).
 We fixed the mid-transit time using simultaneous TESS observations (Feinstein et al.
@@
 We modeled it with a second-order polynomial for the out-of-transit baseline.
-See Appendix C for more discussion on the baseline modelling.
-The planetary transit signal was modelled using batman (Kreidberg 2015).
+See Appendix C for more discussion on the baseline modeling.
+The planetary transit signal was modeled using batman (Kreidberg 2015).
 For the white light curve (1.12-1.6 m), we fit the RECTE parameters and the baseline polynomial coefficients, along with the planet’s transit depth and semi-major axis.
@@
 2024), we were able to significantly reduce the uncertainties on the transit spectrum (Figure 9).
+V1298 Tau b and c have impact parameters (0.34 and 0.46 respectively) consistent within 1 of each other.
+Furthermore, Johnson et al.
+(2022) report a relatively low mutual inclination (019∘), indicating a planar orbital geometry of the V1298 Tau system.
+Therefore, both planets are likely to transit a similar part of the stellar photosphere with limb darkening during their primary transit.
 This highlights the advantage and uniqueness of studying multiple planets in the same system.
-For a discussion on the limb darkening of V1298 Tau see Appendix 4.2.
-We also independently measure the stellar density from the HST white light curves which is discussed in Appendix 4.1.
+For a discussion on the limb darkening of V1298 Tau see Section 4.2.
+We also independently measure the stellar density from the HST white light curves which is discussed in Section 4.1.
 The fit was performed using emcee (Foreman-Mackey et al.
 2013).
-The posterior distribution from these fits is shown in Supplementary Figure 5 and the best-fit transit model and residuals, are shown in Supplementary Figure 4 (upper panel).
+The posterior distribution from these fits is shown in Fig.
+6 and the best-fit transit model and residuals, are shown in Figure 1 (upper panel).
 Spectroscopic light curves were analyzed similarly by dividing the stellar spectra into 7-pixel bins, resulting in 17 channels (identical to V1298 Tau b).
 We fixed the semi-major axis from the white light curve fits and fit the RECTE parameters and baseline for each bin, and the transit depth.
-We fix the wavelength dependent limb-darkening coefficients to values derived in Barat et al.
+We fix the wavelength dependent limb darkening coefficients to values derived in Barat et al.
 (2024).
-The detrended spectroscopic light curves and residuals are shown in Supplementary Figure 6.
-Appendix C Modelling the baseline
-Young stars exhibit quasi-periodic flux modulation due to rotational variability (e.g see Feinstein et al.
-2022; David et al.
+The de-trended spectroscopic light curves and residuals are shown in Fig.
+7.
+Appendix C Modeling the baseline
+Young stars exhibit quasi-periodic flux modulation due to rotational variability (e.g.
+see David et al.
 2019b; Mann et al.
@@
 The transit of planet c analyzed in this work occurred near an inflection point of the stellar baseline, resulting in significant deviation from a linear function.
-We modeled the baseline using linear, quadratic, and cubic polynomial functions for all the spectroscopic channels.
-The transit depths from all three fits were consistent within 1, but quadratic and cubic polynomials were statistically favored over a linear baseline (See Supplementary Figure 7).
+We modelled the baseline using linear, quadratic, and cubic polynomial functions for all the spectroscopic channels.
+The transit depths from all three fits were consistent within 1, but quadratic and cubic polynomials were statistically favored over a linear baseline (See Fig.
+8).
 Higher-order polynomials, however, introduce correlation between the transit depth and baseline, increasing uncertainties and scatter in the transmission spectrum.
@@
 Therefore, we used a quadratic polynomial to model the baseline for the final transmission spectrum.
-Appendix D Atmospheric models with ATMO
-ATMO’s python based retrieval framework was applied and bench-marked in Lewis et al.
-(2020) and Rathcke et al.
-(2021) for HAT-P-41b and WASP-79b, respectively .
-This same retrieval framework is applied to V1298 Tau b and c in this work.
-This framework includes a Nested sampling Bayesian sampler DYNESTY (Speagle 2020) coupled to ATMO.
-We perform chemical equilibrium retrieval with isothermal - profile.
-Chemical equilibrium abundances are computed using Gibbs energy minimization as detailed in Drummond et al.
+Appendix D Theoretical evolutionary models
+The evolutionary models we use in the current study are described in Vazan et al.
 (2022).
-We include H2-H2, H2-He Collision Induced opacity (CIA), as well as H2O, CH4 and CO high temperature molecular opacity in the retrieval.
-The clouds are treated as grey cloud deck between 1 and 50 millibar pressure levels, but with variable cloud opacity governed by the cloud factor used as the variable in the retrieval, as detailed in Goyal et al.
-(2018).
-Appendix E Comparing V1298 Tau b ATMO retrieval with Barat et al. (2024)
-For this study, we performed atmospheric retrieval on V1298 Tau b using the same setup as for planet c, assuming an isothermal atmosphere.
-Our results, indicating sub-solar metallicity and a mass upper limit of 20M⊕, align with Barat et al.
-(2024), who used PetitRadtrans.
-However, our retrieved temperatures are higher (800-900K) compared to their 600-700K.
-The discrepancy arises because Barat et al.
-(2024) used a non-isothermal temperature-pressure (T-P) profile with high interior temperature and vertical mixing to explain the non-detection of methane for this planet.
-Our isothermal model lacks this flexibility, leading to higher temperatures beyond the CH4/CO transition (Fortney et al.
-2020) to eliminate methane from the upper atmosphere.
-Appendix F Haze formation models
-The haze microphysics models were simulated using the models described in Kawashima & Ikoma (2018), following the approach of Kawashima & Ikoma (2019).
-First, using a photochemical model the abundance of gaseous species was calculated.
-The XUV spectrum of V1298 Tau was adopted from Duvvuri et al.
-(2023).
-Haze production rates (or efficiency of haze formation) are uncertain, therefore we assumed a certain fraction of precursor molecules to be dissociated to form haze monomers.
-The haze precursor molecules are assumed to be CH4, HCN and C2H2 in our model.
-We derived vertical density and radius distributions of the haze particles following Kawashima & Ikoma (2018).
-We assumed solar elemental abundance ratio and other parameters such as eddy diffusion coefficient, monomer radius, internal density, and refractive index of haze particles similar to Kreidberg et al.
-(2022).
-We simulated transmission spectra, for V1298 Tau b and c assuming 23 and 20M⊕ respectively with obtained profiles of gas and haze particles.
-In Figure 1 we show the simulated transmission spectra assuming a 1 haze production efficiency.
-Appendix G Theoretical evolutionary models
 Our models start from the time of disk dispersal and calculate the thermal evolution and structural evolution.
@@
 Our models include core erosion and mixing with the envelope.
-The evolutionary models are described in Vazan et al.
-(2022).
 We choose the initial conditions (mass, radius) such that it matches the current mass and radius for the planets.
-We assumed 23M⊕ and 20M⊕ for the current mass estimates of the V1298 Tau b and c respectively.
+We assumed 17 M⊕ (peak of ATMO retrieval posterior distribution) for the current mass estimates of the V1298 Tau c.
+For planet b, we show two models: 10 M⊕ (consistent within 1 to median mass retrieved from ATMO) and 23 M⊕ (3 mass upper limit from Barat et al.
+(2024)).
 We run simulations with both core-envelope and gradually mixed interior models.
@@
 Mass and radius (normalized at current value) evolution models for the V1298 Tau planets are shown in Figure 4.
-Appendix H Effect of stellar activity on transmission spectrum
+Appendix E Effect of stellar activity on transmission spectrum
 We estimate the effect of stellar activity on the transmission spectrum of V1298 Tau c.
@@
 We adopt an extreme spot contrast of 1000 K following spot contrasts derived for a sample of T Tauri stars (Koen 2016).
-We show a comparison between the observed and corrected transmission spectrum in Supplementary Figure 10.
+We show a comparison between the observed and corrected transmission spectrum in Fig.
+11.
 The observed and contamination corrected normalized spectra are consistent within 1 for all spectroscopic channels.
@@
 2021).
-Therefore, the effect of the stellar correction is to change the level of the continuum in this case.
+Therefore, the main effect of the stellar correction is to change the level of the continuum in this case.
 However, the atmospheric features, which are relative to the continuum remain largely unaffected.
-Appendix I Supplementary Figures
-| Planet name | Mass [M⊕] | Atmospheric metallicity [solar] | Semi-major axis [AU] | Stellar metallicity [dex] | Reference |
-| Wasp-18b | 10.200.35 | 1.03 | 0.02 | 0.110.08 | Coulombe et al. (2023) |
-| Wasp-39b | 0.2800.03 | 10 | 0.048 | -0.030.1 | Rustamkulov et al. (2023) |
-| Wasp-77Ab | 1.660.06 | 0.12 | 0.023 | -0.010.1 | August et al. (2023) |
-| Wasp-96b | 0.480.03 | 7.5 | 0.045 | 0.140.19 | Radica et al. (2023) |
-| TOI-270d | 0.0150.001 | 225 | 0.07 | -0.20.1 | Benneke et al. (2024) |
-| Wasp-107b | 0.10.005 | 10 | 0.05 | 0.020.09 | Welbanks et al. (2024) |
-| HD149026b | 0.3580.01 | 20 | 0.0436 | 0.330.1 | Gagnebin et al. (2024) |
-| HD209458b | 0.59029040.01 | 3 | 0.05 | -0.010.05 | Xue et al. (2024) |
-| Wasp-80b | 0.530.03 | 6 | 0.03 | -0.10.1 | Bell et al. (2023) |
-| Wasp-17b | 0.780.2 | 50 | 0.051 | -0.070.1 | Grant et al. (2023) |
-| GJ1214b | 0.0260.001 | 100 | 0.01 | 0.020.1 | Kempton et al. (2023) |
+Appendix F Supplementary Figures
+| Planet name | Mass [MJ] | Atmospheric metallicity [solar] | Semi-major axis [AU] | Stellar metallicity [dex] |
+| WASP-18b 11111111Coulombe et al. (2023) | 10.200.35 | 1.03 | 0.02 | 0.110.08 |
+| WASP-39b 2,3,4,523452,3,4,52 , 3 , 4 , 52,3,4,523452,3,4,52 , 3 , 4 , 5Rustamkulov et al.
+(2023); Feinstein et al.
+(2023); Ahrer et al.
+(2023); Alderson et al.
+(2023) | 0.2800.03 | 10 | 0.048 | -0.030.1 |
+| WASP-77Ab 66666666August et al. (2023) | 1.660.06 | 0.12 | 0.023 | -0.010.1 |
+| WASP-96b 77777777Radica et al. (2023) | 0.480.03 | 0.23 | 0.045 | 0.140.19 |
+| TOI-270d 88888888Benneke et al. (2024) | 0.0150.001 | 225 | 0.07 | -0.20.1 |
+| WASP-107b 99999999Sing et al. (2024) | 0.10.005 | 43 | 0.05 | 0.020.09 |
+| HD149026b 1010101010101010Gagnebin et al. (2024) | 0.3580.01 | 20 | 0.0436 | 0.330.1 |
+| HD209458b 1111111111111111Xue et al. (2024) | 0.59029040.01 | 3 | 0.05 | -0.010.05 |
+| WASP-80b 1212121212121212Bell et al. (2023) | 0.530.03 | 6 | 0.03 | -0.10.1 |
+| WASP-17b 1313131313131313Grant et al. (2023) | 0.780.2 | 50 | 0.051 | -0.070.1 |
+| GJ1214b 1414141414141414Kempton et al. (2023) | 0.0260.001 | 100 | 0.01 | 0.020.1 |